### v1

# Distribution estimation via Flow Matching with Lipschitz guarantees

Distribution estimation via Flow Matching with Lipschitz guarantees
Lea Kunkel [^1] The author thanks Mathias Trabs for inspiring discussions and many helpful comments on this project.
(Karlsruhe Institute of Technology)
Flow Matching, a promising approach in generative modeling, has recently gained popularity. Relying on ordinary differential equations, it offers a simple and flexible alternative to diffusion models, which are currently the state-of-the-art. Despite its empirical success, the mathematical understanding of its statistical power so far is very limited. This is largely due to the sensitivity of theoretical bounds to the Lipschitz constant of the vector field which drives the ODE. In this work, we study the assumptions that lead to controlling this dependency. Based on these results, we derive a convergence rate for the Wasserstein $1$ distance between the estimated distribution and the target distribution which improves previous results in high dimensional setting. This rate applies to certain classes of unbounded distributions and particularly does not require $\log$-concavity.
Keywords: Generative models, continuous normalizing flows, rate of convergence, Wasserstein distance, distribution estimation
MSC 2020: 62E17, 62G07, 68T07
1 Introduction
Generative models aim to learn a mapping $\psi$ that pushes a simple latent variable $Z\sim\mathbb{U}$ towards a target distribution $\mathbb{P}^{*}$ on $\mathbb{R}^{d}$. This enables quick generation through sampling from the latent distribution and applying the learned mapping.
Flow Matching, introduced by Lipman et al. (2023), has recently attracted attention for its simplicity compared to diffusion models (Sohl-Dickstein et al., 2015), which have long been regarded as state-of-the-art. Unlike generative adversarial networks (Goodfellow et al., 2014), which learn $\psi$ directly, Flow Matching learns a time-depended vector field $v_{t}$ that transports mass from $\mathbb{U}$ to $\mathbb{P}^{*}$ via the ODE for $t\in[0,1]$:
$\frac{d}{dt}\psi_{t}(x)=v_{t}\left(\psi_{t}(x)\right),\quad\psi_{0}(x)=x$ (1)
inducing path densities $p_{t}=[\psi_{t}]_{\#}p_{0}$, where $p_{0}$ is the density of the latent distribution $\mathbb{U}$. The vector field $v$ is chosen such that $p_{1}$ approximates the target density.
Earlier, normalizing flows (Tabak and Vanden-Eijnden, 2010; Tabak and Turner, 2013), popularized by Rezende and Mohamed (2015) and Dinh et al. (2015) used discrete invertible mappings with the change-of-variables formula. Continuous normalizing flows (Chen et al., 2018) extended this to continuous mappings using neural ODEs trained via maximum likelihood, which circumvented the calculation of the determinant of the Jacobian. However, CNFs still require significant computational costs due to the necessity of ODE simulations (Grathwohl et al., 2019).
Flow Matching avoids these simulations by using a simple least-squares loss
$\begin{gathered}\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}(\cdot\mid Y)\end{subarray}}\left[\left|\tilde{v}_{t}\left(X_{t}\right)-v_{t}\left(X_{t}\mid Y\right)\right|^{2}\right]\end{gathered}$ (2)
where $v_{t}(\cdot|Y)$ and $p_{t}(\cdot|Y)$ are known by design. In some settings, this resembles the score matching objective (Hyvärinen, 2005; Vincent, 2011) used in diffusion models (Song et al., 2020). Given $n$ i.i.d. observations from $\mathbb{P}^{*}$, minimizing the empirical counterpart of (2) over a set of functions $\mathcal{M}$ and solving the ODE (1) leads to the estimator $\hat{\psi}$. The statistical question is how well $\mathbb{P}^{\psi_{1}(Z)}$ approximates $\mathbb{P}^{*}$ depending on the number of observations $n$ and the function class $\mathcal{M}$.
Algorithms based on Flow Matching have been successfully applied in a range of domains, including text-to-speech (Guo et al., 2024), text-to-image generation (Yang et al., 2024; Esser et al., 2024), molecular and protein structure design (Dunn and Koes, 2024; Bose et al., 2024), and surrogate modeling in high-energy physics (Bieringer et al., 2024). The method has also been extended to various theoretical settings: Atanackovic et al. (2024) adapted it to interacting particle systems, Gat et al. (2024) explored discrete spaces, Chen and Lipman (2024) generalized it to Riemannian manifolds, and Kerrigan et al. (2023) extended it to function spaces.
The analysis of Flow Matching is more fragile compared to Diffusions, as Girsanov’s theorem does not apply, thus making the strategy used in Chen et al. (2023c) inapplicable. Using Grönwall’s Lemma, the natural and standard approach for stability estimates of ODEs, results in exponential dependency on the Lipschitz constant of the approximated vector field. Consequently, the statistical results so far are rather limited: Fukumizu et al. (2024) use a slight adaptation and introduce stopping times to transfer some results from diffusion models but they do not circumvent the growth of the Lipschitz constant, which causes difficulties in their proof. Gao et al. (2024b) use the same adaptation and obtain rates in the Wasserstein 2 distance for certain types of unknown distributions. Stéphanovitch et al. (2025) study deterministic ODE sampling, but, due to their general setting, examine the infinite-time setting of Diffusions. Kunkel and Trabs (2025b) showed that Flow Matching attains minimax optimal rates, but they use an over parameterized setting which does not align with practical implementations. Due to the excess of parameters, they are able to compensate the rising Lipschitz constant.
In this work, we first focus on a detailed study of the Lipschitz constant of the intrinsic "true" vector field of Flow Matching. By providing upper and lower bounds, we demonstrate that the choice of the variance function and the behavior of covariance of the reweighted unknown distribution is crucial. This function governs the level of smoothing depending on $t\in[0,1]$. While current results focus on fixed choices of variance functions, we study the Lipschitz constant in terms of this function. We derive assumptions on the unknown distribution that lead to control over the Lipschitz constant for various types of variance functions. Our general assumptions are validated for a class of distributions with unbounded support considered by Stéphanovitch et al. (2025).
Subsequently, we employ Bernstein-type bounds to derive an oracle inequality for estimating the vector field. However, we carefully adapt these classical methods to address the theoretical challenges of Flow Matching. When bounding the error of the Flow Matching distribution estimator in the Wasserstein $1$ distance, standard approaches from the analysis of diffusion models usually increase the negative impact of the dimension on the rate. This issue can be circumvented up to a small error by exploiting the smoothness of the vector field. Using this approach allows us to build on existing results on network approximation. Our focus is on feedforward rectified linear unit (ReLU) networks, hence we do not require smooth networks for our results. ReLU nets with logarithmically increasing depth and a polynomial number of non-zero weights suffice to achieve improved convergence rates. Additionally, we exploit the density path of the true vector field to take advantage of the smoothness of the unknown distribution. Thus, our rates are faster in high dimensions than those previously obtained in Gao et al. (2024b), while allowing for considerably smaller networks than in Kunkel and Trabs (2025b).
Further Related work
Approaches related to flows between two possibly unknown distributions, $\mathbb{P}$ and $\mathbb{Q}$, are studied by Tong et al. (2024), Liu et al. (2022) and Albergo and Vanden-Eijnden (2022). Tong et al. (2024) also generalized and unified these methods. Benton et al. (2024) studied flow matching, excluding approximation errors, by imposing assumptions on the covariance that provided global Lipschitz bounds on the vector field. The work by Gong et al. (2025) examines the properties of ReLU networks used to approximate a vector field corresponding to higher-order trajectories.
As Lipman et al. (2023) noted, Flow Matching is closely connected to diffusion models. For a comprehensive survey of generative diffusion models, refer to the work of Cao et al. (2024). Even when the model is not built to be consistent with diffusion models, approximating a score function is similar to approximating the vector field in Flow Matching. The statistical analysis of score matching and diffusion models are an active area of research, see for instance Chen et al. (2023a), Chen et al. (2023b), Chen et al. (2023c), Oko et al. (2023), Tang and Yang (2024), Azangulov et al. (2024), Zhang et al. (2024) , Yakovlev and Puchkin (2025), Stéphanovitch et al. (2025).
Outline
Section˜2 introduces the problem setting mathematically and recalls all the definitions needed to define the conditional Flow Matching objective from Lipman et al. (2023). Section˜3 then discusses the use of the Wasserstein distance and shows a classical first error decomposition using the aforementioned Grönwall’s lemma. After that, we introduce the model based on the choice $\mathbb{U}=\mathcal{N}(0,I_{d})$, which will be the focus of the subsequent work. In Section˜4 we establish upper and lower bounds of the Lipschitz constant. We then show that bounding the Lipschitz constant requires careful investigation of the covariance of the reweighted unknown distribution. Next, we demonstrate that, under certain assumptions, we can construct Flow Matching such that the Lipschitz constant is bounded. Following that, we present a class of densities that satisfies our assumptions. In Section˜5 we derive an oracle inequality and finally obtain a rate of convergence using feedforward ReLU networks. All proofs are deferred to Section˜7.
1.1 Some notations
We use the notation $X\sim p$ as shorthand for $X\sim\mathbb{P}$ when the distribution $\mathbb{P}$ of the random variable $X$ admits a density $p$ with respect to the Lebesgue measure.
The symbol $\lesssim$ denotes an inequality of the form $a\leq c\cdot b$, where $a,b\in\mathbb{R}$ and $c$ is a constant independent of the sample size $n$. In the same setting, the symbol $a\sim b$ denotes equality up to a constant. For $d\in\mathbb{N}$, we equip $\mathbb{R}^{d}$ with the standard Euclidean norm $|\cdot|$. For a matrix $A\in\mathbb{R}^{d\times d}$, we denote the spectral norm by $\|A\|$. Given a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$, where $\Omega\subset$ $\mathbb{R}^{d^{\prime}}$, the supremum norm and the $L_{1}$-norm are defined by
$\|f\|_{\infty,\Omega}\coloneqq\sup_{x\in\Omega}|f(x)|\quad\text{and}\quad\|f\|_{1,\Omega}\coloneqq\int_{\Omega}|f(x)|\;\mathrm{d}x.$
When $\Omega=\mathbb{R}^{d^{\prime}}$, we abbreviate $\|f\|_{\infty,\mathbb{R}^{d^{\prime}}}$ as $\|f\|_{\infty}$, and similarly $\|f\|_{1,\mathbb{R}^{d^{\prime}}}=\|f\|_{1}$. For $d^{\prime\prime}=1$ and $\alpha\in(0,1]$, the Besov space $B_{1,\infty}^{\alpha}(\Omega)$ is defined as
$B_{1,\infty}^{\alpha}(\Omega)\coloneqq\big{\{}f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}<\infty\big{\}},$
where the modulus of smoothness is given by
$\omega_{1}(f,t)_{1}\coloneqq\sup_{0<|h|\leq t}\int|f(x)-f(x+h)|\;\mathrm{d}x,\quad t>0,x\in\Omega.$
If $x+h\notin\Omega$, the integrand is taken to be zero. The norm on $B_{1,\infty}^{\alpha}(\Omega)$ is defined by
$\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}.$
Lastly, $C^{k}(\Omega)$ denotes the set of functions whose coordinate components are $k$-times continuously differentiable with bounded derivatives (in the supremum norm). For a multi-index $k\in\mathbb{N}_{0}^{d}$ with $|k|=$ $\sum_{i=1}^{d}k_{i}$, we write
$D^{k}=\frac{\partial^{|k|}}{\partial x_{1}^{k_{1}}\cdots\partial x_{d}^{k_{d}}},$
where $\frac{\partial}{\partial x_{i}}$ denotes the weak partial derivative. For a function $f\colon\Omega\rightarrow\mathbb{R}$ we denote for $\beta\in\mathbb{N}$
$\|f\|_{C^{\beta}}\coloneqq\max_{k\colon|k|\leq\beta}\|D^{k}f\|_{\infty}.$
For the Jacobian of a vector valued function with respect to the vector $x$ we write $D_{x}$. The covering number of a set $\mathcal{A}$ with respect to the norm $\|\cdot\|$ and the $\tau$-covering is denoted by $\mathcal{N}(\tau,\mathcal{A},\|\cdot\|)$.
2 The Flow Matching objective
Suppose we observe an i.i.d. sample $X_{1}^{*},\ldots,X_{n}^{*}$ drawn from an unknown distribution $\mathbb{P}^{*}$. Further assume that $\mathbb{P}^{*}$ admits a density $p^{*}$ with respect to the Lebesgue measure. The aim of generative modeling is to mimic the distribution $\mathbb{P}^{*}$.
To this end, consider a time-dependent vector field $v\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, and define the flow $\psi\colon[0,1]\times$ $\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as the solution to the ODE:
$\frac{d}{dt}\psi_{t}(x)=v_{t}(\psi_{t}(x)),\quad\psi_{0}(x)=x.$ (3)
Given a fixed latent distribution $\mathbb{U}$ with Lebesgue density $p_{0}$, the vector field $v$ induces a path of probability densities $p_{t}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$, defined via the push-forward:
$p_{t}=\left[\psi_{t}\right]_{\#}p_{0},\quad\text{ meaning }\quad\psi_{t}(Z)\sim p_{t}\text{ for }Z\sim p_{0},$
with $\int p_{t}(x)\mathrm{d}x=1$ for all $t\in[0,1]$.
Using a possibly parameterized function $\tilde{v}$ to approximate a given vector field $v$ results in the Flow Matching objective
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|\tilde{v}_{t}(X_{t})-v_{t}(X_{t})|^{2}\big{]}$ (4)
by Lipman et al. (2023). Conversely, in certain cases, a vector field can be determined to match a given density path. Lipman et al. (2023) constructed probability paths resulting from a convolution,
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y,$ (5)
where $p_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a conditional probability path induced through some vector field $v_{t}(\cdot|y)\colon\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for $y\in\mathbb{R}^{d}$. It can be shown that the vector field inducing (5) is of the form
$v_{t}(x)=\int v_{t}(x|y)\frac{p_{t}(x|y)p^{*}(y)}{p_{t}(x)}dy.$ (6)
In this setting, Lipman et al. (2023) prove that minimizing the Flow Matching objective (4) over a class of parameterized functions $\mathcal{M}$ is equivalent to minimizing the conditional Flow Matching objective given by
$\Psi(\tilde{v})\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}.$ (7)
As the following result shows, we can calculate the difference between (4) and (7) explicitly.
Lemma 2.1.
Let $p_{t}(x)>0$ for all $x\in\mathbb{R}^{d}$. In the above setting, it holds that for every measurable function $\tilde{v}$ and $v_{t}$ from (6)
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}-\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}[|v_{t}(X_{t})-v_{t}(X_{t}|Y)|^{2}].$
In practice, $p^{*}$ is unknown and therefore the empirical counterpart of (7),
$\tilde{\Psi}(\tilde{v})\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ X_{t}\sim p_{t}\left(\cdot|Y_{i}\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y_{i}\right)\right|^{2}\big{]}$ (8)
is minimized. Kunkel and Trabs (2025b) show that there is also an equivalence of gradients in the empirical setting. Let $\mathcal{M}$ be a set of parameterized functions that map $\mathbb{R}^{d}$ into $\mathbb{R}^{d}$. From minimizing (8) in $\tilde{v}\in\mathcal{M}$, we obtain an argument $\hat{v}$. Solving the ODE (3) using $\hat{v}$, we obtain a flow $\hat{\psi}_{t}$, i.e. $\hat{\psi}$ is given by
[…]
We use this flow to push forward the known, latent distribution to time $t=1$. In accordance with the goal of generative modeling, the distribution of this pushforward should mimic $\mathbb{P}^{*}.$
3 Wasserstein distance in Flow Matching
The underlying objective of this work is to evaluate the ability of Flow Matching to mimic the unknown distribution $\mathbb{P}^{*}$ depending on the sample size $n$. For this purpose, it is necessary to measure the gap between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (9). We adopt the Wasserstein-$1$ metric for evaluation.
On the normed vector space $(\mathbb{R}^{d},|\cdot|)$ the Wasserstein-$1$ distance between two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{d}$ is given by
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})$ $\coloneqq\inf_{\pi\in\Pi(\mathbb{P},\mathbb{Q})}\int_{\mathcal{X}}|x-y|\;\mathrm{d}\pi(x,y),$
where $\Pi(\mathbb{P},\mathbb{Q})$ denotes the set of couplings with marginals $\mathbb{P}$ and $\mathbb{Q}$. By Villani (2008, Theorem 5.10(i)), this definition admits the following dual formulation:
$\mathsf{W}_{1}(\mathbb{P},\mathbb{Q})=\sup_{W\in\operatorname{Lip}(1)}\mathbb{E}_{\begin{subarray}{c}X\sim\mathbb{P}\\ Y\sim\mathbb{Q}\end{subarray}}[W(X)-W(Z)].$ (10)
On the space of probability measures with finite first moment, the Wasserstein-$1$ distance characterizes weak convergence. A detailed discussion of its advantages over other metrics with the same characteristic can be found in (Villani, 2008, p. 98 f.). Due to its geometric interpretability and robustness in high-dimensional, manifold-structured data settings, the Wasserstein-$1$ metric has commonly been used to evaluate generative models. Examples include: Schreuder et al. (2021), Liang (2021), Stéphanovitch et al. (2023) or Kunkel and Trabs (2025a).
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and $\mathbb{P}^{\psi_{1}(Z)}$, where $\psi$ is the solution of the ODE (3) using (6) as the vector field. In order to bound $\mathsf{W}_{1}(\mathbb{P}^{\hat{\psi}_{1}(Z)}\mathbb{P}^{\psi_{1}(Z)})$ in terms of the difference between the underlying vector fields, we use Grönwall’s Lemma as the classical tool to control ODE stability. This leads to the typical exponential bound on the Lipschitz constant in space. The proof follows the along lines of (Albergo and Vanden-Eijnden, 2022, Proposition 3).
Theorem 3.1.
(Kunkel and Trabs, 2025b, Theorem 3.1) Assume that all functions in $\mathcal{M}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\Gamma_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{M}$ and $v$ from (6)
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}(Z)})+\sqrt{2e}e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}\Big{(}\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}\Big{)}^{1/2}.$ (11)
The first term in (11) is determined solely by the conditional vector fields $v_{t}(\cdot|y)$and is therefore independent of the approximation class $\mathcal{M}$. To bound this term, we impose the following assumptions on the densities of the unknown distribution $\mathbb{P}^{*}$ and the latent distribution $\mathbb{U}$.
Assumption 3.2.
We assume that
  1. 1.
    $d\geq 2,p^{*}\in B_{1,\infty}^{\alpha},\alpha\in(0,1]$.
  2. 2.
    $\int yp_{0}(y)\;\mathrm{d}y=0,\quad\text{and}\quad\int|y|^{1+\alpha}p_{0}(y)\;\mathrm{d}y<\infty.$
Under ˜3.2, we know that
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}(Z)})\lesssim\sigma_{\min}^{1+\alpha},$ (12)
see proof of (Kunkel and Trabs, 2025b, Theorem 3.3), where the first part of the proof does not use the assumption that the support of $\mathbb{P}^{*}$ is compact. Note that the constant in (12) depends only on $\|p^{*}\|_{B^{\alpha}_{1,\infty}}$ and $d$.
The second term in (11) is connected to the suitability of the set $\mathcal{M}$ to approximate $v$. This depends on the specific structure of $v$. The factor $\exp(\int_{0}^{1}2\Gamma_{t}\;\mathrm{d}t)$ results from the use of Grönwall’s inequality and is one of the key difficulties in the analysis of Flow Matching. In view of Equation˜4, any Lipschitz constraint on the class $\mathcal{M}$ should be aligned with the Lipschitz constant of $v.$ Looking at (6), this Lipschitz constant depends on the construction of $v_{t}(\cdot|\cdot)$ and $p_{t}(\cdot|\cdot)$ and the unknown density $p^{*}$.
In the next section, we are going to study this Lipschitz constant. We are going to focus on one specific probability path and assume the following:
Assumption 3.3.
We assume that
$p_{t}(x|y)\propto\exp\Big{(}-\frac{|x-\mu_{t}(y)|^{2}}{2\sigma^{2}_{t}}\Big{)},$
where the mean shift $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and the variance function $\sigma\colon[0,1]\rightarrow\mathbb{R}_{>0}$ are smooth, monotone functions such that
$\mu_{0}(y)=0,$ $\quad\mu_{1}(y)=y,$
$\sigma_{0}=1,$ $\quad\sigma_{1}=\sigma_{\min},$
for a $\sigma_{\min}>0$.
˜3.3 means in particular that the latent distribution is $\mathcal{N}(0,I_{d})$ The monotonicity assumption guarantees that there is no unnecessary movement of mass. ˜3.3 coincides with the construction in Lipman et al. (2023, Section 4). Very similar generative models based on probability Flows have been studied by Gao et al. (2024b, a). Additionally, the use of $\mathcal{N}(0,I_{d})$ for the latent distribution is pervasive in generative modeling, e.g. in Diffusions building up on Sohl-Dickstein et al. (2015).
4 Lipschitz constant of the vector field
Under ˜3.3 Lipman et al. (2023), show that the corresponding vector field that generates $p_{t}(\cdot|y)$ for every $y\in\mathbb{R}^{d}$ is given by
$v_{t}(x|y)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-\mu_{t}(y))+\mu_{t}^{\prime}(y),$ (13)
where $\sigma^{\prime}_{t}$ and $\mu^{\prime}_{t}$ are the derivatives in time. We conclude from this result that the definition of the variance function $\sigma_{t}$ is critical for bounding the Lipschitz constant, while only very special choices of the mean shift $\mu_{t}$ have an influence. Therefore, we will focus on polynomial choices of $\mu_{t}$, which generalize the choice of Lipman et al. (2023).
Assumption 4.1.
We consider the following choices of $\mu\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ for $\gamma>1$:
$\mu_{t}(y)=t^{\gamma}y.$
A natural first idea is to aim for a global Lipschitz constant, this corresponds to the assumption that $v$ has a bounded Lipschitz constant. The next result shows that such a constant cannot exist in this setting. We can upper and lower bound the Lipschitz constant of $v$:
Theorem 4.2.
Let
$B^{t}_{i,j}\coloneqq\Big{\|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\Big{\|}_{\infty},$
where $Y^{x,t}$ is a random variable with density $q\propto p_{t}(x|\cdot)p^{*}(\cdot)$ for $x\in\mathbb{R}^{d}$. Assume $B_{i,j}$ exists for all $i,j\in\{1,..,d\}$ and define $B\coloneqq(B_{ij})_{i,j=1,...,d}$.
For the Lipschitz constant in space $\Gamma_{t}$ of $v_{t}$ on $\mathbb{R}^{d}$, we have that
$\max_{ij}B_{i,j}\leq\Gamma_{t}\leq d\max_{ij}B_{i,j}.$
The upper and lower bound in Theorem˜4.2 depend on the choice of $\sigma_{t}$. We want to choose $\sigma_{t}$ such that $\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}$ does not grow too fast for small $\sigma_{t}$ or large $|\sigma_{t}^{\prime}|$. The next result shows that it is not possible to choose a function $\sigma_{t}$ such that the absolute value of the quotient is smaller than $\log(\sigma_{\min}^{-1})$ for all $t\in[0,1]$ or that the integral over $\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}$ remains small.
Lemma 4.3.
  1. 1.
    There is a $t^{*}\in[0,1]$ such that
    $\frac{\sigma_{t^{*}}^{\prime}}{\sigma_{t^{*}}}=\log(\sigma_{\min}),$
    for any choice of $\sigma_{t}$ satisfying ˜$3.$
  2. 2.
    For any choice of $\sigma_{t}$ satisfying ˜3.3,
    $\int_{0}^{1}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t=\log(\sigma_{\min}^{-1}).$
We note that Lemma˜4.3 does not depend on the specific probability path, it holds for all quotients of smooth choices of $\sigma_{t}$. Furthermore, similar terms arise in all cases, where the chosen probability path implies vector fields of the form (13). The first part of Lemma˜4.3 reveals that a global Lipschitz bound on $v_{t}$ independent of $\sigma_{\min}$ and the covariance term is not feasible in this setting. To compensate for $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min})$ in the case of $i=j$, the second term must have the same absolute value. However, this prohibits a smaller bound when $i\neq j$. The second part of Lemma˜4.3 implies that, even the case of $\operatorname{Cov}(Y^{x,t})_{ij}\lesssim\sigma_{t}^{2}$ for all $i,j$ will lead to a logarithmic dependency on $\sigma_{\min}^{-1}$, which ultimately influences the rate.
To control the Lipschitz constant, we thus need to assume that $p^{*}$ is such that the covariance term in Theorem˜4.2 decays in a controlled order for $\sigma_{t}\rightarrow\sigma_{\min}$ for $i=j$ and the same decay of the off-diagonal elements is fast enough.
Assumption 4.4.
Assume there is a $t^{*}\in\big{[}\frac{1}{2^{1/\gamma}},1\big{)}$ independent from $\sigma_{\min}$ such that for all $x\in\mathbb{R}^{d}$
$\mathrm{(I)}$ $\operatorname{Cov}(Y^{x,t})_{ij}$ $\lesssim\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{3},$ $i$ $\neq j,$ $t>t^{*},$
$\mathrm{(II)}$ $\operatorname{Var}(Y^{x,t}_{i})$ $=\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}\big{(}1+O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}),$ f $\text{or all }i,$ $t>t^{*},$
$\mathrm{(III)}$ $\operatorname{Cov}(Y^{x,t})_{ij}$ $\leq C,$ f $\text{or all }i,j,$ $t\leq t^{*},$
where $Y^{x,t}$ is a random variable with density $q\propto p_{t}(x|\cdot)p^{*}(\cdot)$ and $\kappa\in\mathbb{R}_{\geq 1}$ and $C$ are fixed constants.
Under these assumptions, any choice of $\sigma_{t}$ satisfying ˜3.3 will lead to a bounded Lipschitz constant.
Theorem 4.5.
Grant ˜3.3, ˜4.1 and ˜4.4 with fixed parameters $\gamma,\kappa$ and $t^{*}$. Then there is a constant $C$ such that
$\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t\leq C.$
In order to allow for comparison with other works, we are going to show that the very abstract ˜4.4 is satisfied if the unknown distribution is of the form
$p^{*}(x)\propto\exp\big{(}-\frac{|x|^{2}}{2}-a(x)\big{)},\quad\|a\|_{C^{2}}=L<\infty.$ (14)
This setting was used in Stéphanovitch (2024) to construct Lipschitz continuous pushforward maps in diffusion models.
This setting specifically allows for unbounded distribution that are not $\log$-concave (i.e the logarithmic density is concave) and thus the Brascamp-Lieb inequality (Brascamp and Lieb, 1976, Theorem 4.1) cannot be applied directly. The $\log$-concavity assumption has been popular in the analysis of Diffusion models and Flow Matching, see for example Gao and Zhu (2024); Bruno et al. (2023); Gao et al. (2025, 2024b). First, we note that by Lemma˜7.11 the first assumption in ˜3.2 is satisfied for densities of the form (14) with $\alpha=1$. For simplicity, we choose $\gamma=1$, but the result extends easily to other $\gamma$.
Theorem 4.6.
Set $\gamma=1$. Assume that $\mathbb{P}^{*}$ is of the form (14). Then ˜4.4 is fulfilled with
$C=e^{2L},\quad\kappa=1,\quad\text{ and }\quad t^{*}\quad\text{ such that }\quad\sigma_{t^{*}}=c(L,d),$
where $c(L,d)\in(0,1)$ is a constant that depends only on $L$ and $d$.
The uniform covariance bound is proven using the fact that we can control the effect of bounded perturbations of Gaussians on the variance, which follows from the Holley-Strooke perturbation principle. The bound on the diagonal entries employs proof techniques from the Cramér-Rao lower bound. The bound on the off-diagonal elements is developed from the Brascamp-Lieb type covariance estimate of Menz (2014).
The proof of Theorem˜4.6 reveals that for $t>t^{*}$ large enough and $i\neq j$, the covariance decay is of order $\sigma_{t}^{4}$ and $\kappa=1$. This suggests that the class of distributions for which Theorem˜4.5 applies is significantly larger.
5 Rate of convergence
In order to evaluate the performance of Flow Matching, we are going to derive a rate of convergence using neural networks for the set $\mathcal{M}.$ Since $\tilde{v}$ is chosen such that it minimizes (8) instead of (4), we cannot use the minimization property in Theorem˜3.1 directly. The next theorem provides a more detailed decomposition of the error in the form of a classical oracle inequality. In order to merge the results later, we choose $\mu_{t}$ such that the results of Section˜4 apply. For the variance function $\sigma_{t}$, the analysis in Section˜4 reveals that we are not restricted to the linear case. Thus, we can choose a variance function that is suited for the application of a Bernstein-type inequality. Since Bernstein-type inequalities rely on absolute value bounds, we are going to choose $\sigma_{t}$ such that the absolute value of (13) is as small as possible. Our choice follows from Lemma˜4.3 and the solution of the ODE
$\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{{\min}}),\quad\sigma_{0}=1.$
Assumption 5.1.
Assume that
$\mu_{t}(y)=ty\quad\text{and}\quad\sigma_{t}=(\sigma_{\min})^{t}.$
First, we choose $\mathcal{M}$ as a set of continuous measurable functions $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that for the $j$-th component function $\tilde{v}^{j}$ of $\tilde{v}$
$|\tilde{v}^{j}|_{\infty}\leq e^{2L}\log(n)^{3}+(1+e^{2L})\log(n)^{2}+\log(n)+1.$ (15)
We are going to justify this bound in the proof of the following theorem.
Theorem 5.2.
Let $p^{*}$ be of the form (14) and grant ˜3.3 and ˜5.1. Assume that $\log(\sigma_{\min})\sim\log(n)$. Then we have for every $a\in(0,1],\tau\in\mathbb{R}_{>0},b>1$ and $n$ big enough with probability of $1-\frac{1}{n}$ that
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\hat{v}_{t}(X_{t})|^{2}\big{]}$ $\lesssim\inf_{\tilde{v}\in\mathcal{M}}\int\int|\tilde{v}_{t}(x)-v_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t+n^{-\frac{1}{2}}$ (16)
$\quad+\frac{a^{-1}\log(n)^{7}}{n}\log\big{(}2\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})\big{)}+(2+a)\tau+a\log(n)^{6},$
where $\tilde{v}\in\mathcal{M}$ is chosen such that it is arbitrarily close to realizing the infimum in (16).
The proof of Theorem˜5.2 uses the Bernstein-type concentration inequality of Chen et al. (2023b), which is commonly used in the analysis of diffusion models, see for example Yakovlev and Puchkin (2025). Theorem˜5.2 shows that a careful choice of the variance shift can allow for to a logarithmic dependency of the oracle bounds on $\sigma_{\min}^{-1}$ instead of a linear dependency.
In practice, the parameterized function $\hat{v}$ is typically represented by a neural network. To analyze the theoretical performance in this setting, we connect Theorem˜5.2 with known approximation results for neural networks. Specifically, we consider the following fully connected feedforward networks with ReLU activation functions, which have been extensively studied (see, e.g., Yarotsky (2017), Gühring et al. (2020), Kohler and Langer (2021), Schmidt-Hieber (2020), Suzuki (2019)).
The Rectified Linear Unit (ReLU) activation is defined as $\phi\colon\mathbb{R}\rightarrow\mathbb{R}$, with
$\phi(x)\coloneqq\max(0,x).$
For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, the shifted ReLU activation $\phi_{v}\colon\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is defined componentwise by
$\phi_{v}(x)\coloneqq\left(\phi\left(x_{1}-v_{1}\right),\ldots,\phi\left(x_{p}-v_{p}\right)\right),\quad x=\left(x_{1},\ldots,x_{p}\right)\in\mathbb{R}^{p}.$
A neural network with $L\in\mathbb{N}$ hidden layers and architecture $\mathcal{A}=\left(p_{0},p_{1},\ldots,p_{L+1}\right)\in\mathbb{N}^{L+2}$ is a function
$f\colon\mathbb{R}^{p_{0}}\rightarrow\mathbb{R}^{p_{L+1}},\quad f(x)\coloneqq W_{L}\circ\phi_{v_{L}}\circ W_{L-1}\circ\phi_{v_{L-1}}\circ\cdots\circ W_{1}\circ\phi_{v_{1}}\circ W_{0}\circ x,$ (17)
where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ are weight matrices and $v_{i}\in\mathbb{R}^{p_{i}}$ are shift vectors.
Using networks of the form (17) clipped at (15) for the set $\mathcal{M}$, the next result combines the findings of Section˜4, (12) and Theorem˜5.2 with approximation theory of Gühring et al. (2020) that allows for simultaneous approximation of a function and its derivative. To facilitate the transfer of results, we refrain from inserting $\alpha=1$, which is the applicable smoothness in this setting as shown in Lemma˜7.11.
Theorem 5.3.
Let $p^{*}$ be of the form (14) and grant ˜3.3 and ˜5.1. Then for $n$ large enough with probability of $1-\frac{1}{n}$ for fixed $\eta>0$
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\lesssim\operatorname{polylog}(n)n^{-\frac{1+\alpha}{d+4\alpha+5+\eta}},$
where $\hat{\psi}$ is the solution of an ODE whose vector field is in $\mathcal{M}$ given by a ReLU neural network, with no more than $c\cdot\log\left(n\right)$ layers, $c\cdot n^{c(d,\alpha,\eta)}\cdot\log^{2}\left(n\right)$ nonzero weights, where $c$ and $c(d,\alpha,\eta)$ is a constant independent of $n$.
[…]
Compared to Kunkel and Trabs (2025b, Theorem 3.6), our result applies to networks with logarithmically growing depth and polynomially growing numbers of non-zero weights. On the other hand, the rate deteriorates slightly. However, the two results cannot be compared directly. Kunkel and Trabs (2025b, Theorem 3.6) assume that the support of $p^{*}$ is compact, whereas Theorem˜5.3 assumes full support on $\mathbb{R}^{d}$. The evaluation metric differs from that in Gao et al. (2024b), who use the Wasserstein $2$ metric. According to Villani (2008, Remark 6.5), their rate $n^{-\frac{1}{d+5}}$ is also valid for the Wasserstein $1$ metric. For $\alpha\neq 0$ and large $d$, the rate obtained in Theorem˜5.3 is faster. However, this is just a rough comparison, as the unknown distributions studied are different. Furthermore, they use a linear variance shift and an early stopping approach, which hinders direct comparison even more.
6 Limitations and outlook
Our analysis provides a theoretical explanation for the empirical success of Flow Matching and aligns more closely with real-world scenarios than previous results.
Section˜4 paves the way for further research into broader classes of distributions whose vector fields have a bounded Lipschitz constant. The class of functions of form (14) served as a toy example and is an interesting starting point for generalization. Conversely, the lower bound on the Lipschitz constant could be used to identify distributions that cannot be mimicked with a "good" rate of convergence. Another promising direction is the introduction and consequences of an artificial Lipschitz control of the vector field. This influences the equivalence of gradients between (4) and (7) as well as their empirical counterparts. Controlling this effect could extend convergence results to larger classes of unknown distributions.
Finally, using specifically tailored approximation results or more sophisticated neural networks could improve the convergence rate further.
7 Proofs
7.1 Proofs of Section˜2
Proof of Lemma˜2.1.
From Lipman et al. (2023), we know that there is a constant $C\in\mathbb{R}$ independent of $\tilde{v}$ such that
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}+C.$
Setting $\tilde{v}_{t}=v_{t}$, we obtain
$0=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|v_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}+C\quad\Longleftrightarrow\quad C=-\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|v_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}.$
This concludes the proof. ∎
7.2 Proofs of Section˜4
Proof of Theorem˜4.2.
First we show that we can calculate the Jacobian of $v$ explicitly. The proof and all subsequent proofs of auxiliary results are deferred to Section˜7.4.
Lemma 7.1.
Fix $t\in[0,1].$ The Jacobian with respect to $x$ of $v_{t}$ is given by
$D_{x}v_{t}(x)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}I_{d}+\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t}).$ (18)
The matrix $(B_{ij})_{i,j=1,...,d}$ consists of the component wise supremum norm of $D_{x}v_{t}$.
By assumption $\operatorname{Cov}(\operatorname{Cov}(Y^{x,t})_{ji})$ is bounded for all $x$. For the upper bound we use that by the mean value theorem, there exists an $\xi\in\mathbb{R}^{d}$ such that for the $i$-th coordinate function $v_{t}^{j}$
$|v^{j}_{t}(x)-v^{j}_{t}(y)|=\langle\nabla v^{j}_{t}(\xi),x-y\rangle\leq|v^{j}_{t}(\xi)||x-y|\leq\sqrt{d}|x-y|\max_{i\in\{1,...,d\}}\big{\|}\frac{\partial}{\partial x_{i}}v^{j}_{t}\big{\|}_{\infty}.$
Then
$|v_{t}(x)-v_{t}(y)|=\Bigg{|}\left(\begin{array}[]{c}v^{1}_{t}(x)-v^{1}_{t}(y)\\ \vdots\\ v^{d}_{t}(x)-v^{d}_{t}(y)\end{array}\right)\Bigg{|}\leq d|x-y|\max_{j\in\{1,...d\}}\max_{i\in\{1,...,d\}}\big{\|}\frac{\partial}{\partial x_{i}}v^{j}_{t}\big{\|}_{\infty}.$
Therefore
$|v_{t}(x)-v_{t}(y)|\leq d|x-y|\max_{ij}\big{\|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\Big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\Big{\|}_{\infty}=d|x-y|\max_{ij}B_{i,j}.$
Hence $v_{t}$ is Lipschitz continuous. For the lower bound we can use that by a Taylor expansion for $h,a\in\mathbb{R}^{d}$
$v_{t}(a+h)=v_{t}(a)+D_{x}v_{t}(a)h+r(h),$ (19)
with
$\lim_{|h|\rightarrow 0}\frac{|r(h)|}{|h|}=0.$
Let the smallest Lipschitz constant of $v_{t}$ be $\Gamma_{t}$. Using (19) we can conclude
$|D_{x}v_{t}(a)h|=|v_{t}(a+h)-v_{t}(a)-r(h)|$ $\leq|v_{t}(a+h)-v_{t}(a)|+|r(h)|$
$\leq\Gamma_{t}|h|+|r(h)|.$
Now
$\|D_{x}v_{t}(a)\|=\underset{|h|\rightarrow 0}{\lim\operatorname{sup}}\,\frac{|D_{x}v_{t}(a)h|}{|h|}\leq\underset{|h|\rightarrow 0}{\lim\operatorname{sup}}\,\Gamma_{t}+\frac{|r(h)|}{|h|}=\Gamma_{t}.$
Let $v^{i}_{t}$ be the $i$-th component function of $v_{t}$. Then for every $i\in\{1,...,d\}$ and every $j\in\{1,...,d\}$
$\sup_{a\in\mathbb{R}^{d}}\|D_{x}v_{t}(a)\|=\sup_{a\in\mathbb{R}^{d}}\sup_{|w|=1}|D_{x}v_{t}(a)w|$ $\geq\sup_{a\in\mathbb{R}^{d}}\sup_{|w|=1}|e_{i}^{\top}D_{x}v_{t}(a)w|=\sup_{a\in\mathbb{R}^{d}}\sup_{|w|=1}|\langle\nabla v^{i}_{t}(a),w\rangle|.$
As the dual norm of the euclidean norm is the euclidean norm,
$\sup_{a\in\mathbb{R}^{d}}\sup_{|w|=1}|\langle\nabla v^{i}_{t}(a),w\rangle|=\sup_{a\in\mathbb{R}^{d}}|\nabla v^{i}_{t}(a)|\geq\sup_{a\in\mathbb{R}^{d}}|D_{x}v_{t}(a)_{ij}|.$
Since $i,j$ were arbitrary and the entries of the Jacobian are of the form (18), we obtain the result. ∎
Proof of Lemma˜4.3.
  1. 1.
    We begin by setting
    $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=h_{t},$
    where $h_{t}$ is a continuous function on $[0,1].$ By separations of variables, all of the solutions of this ODE are of the form
    $\sigma_{t}=ce^{H_{t}},$
    where $H_{t}$ is an anti-derivative of $h_{t}$ and $c\in\mathbb{R}$. we use $\sigma_{0}=1$ as initial condition, which leads to
    $1=ce^{H_{0}}\quad\Longleftrightarrow\quad c=\frac{1}{e^{H_{0}}}.$
    To assure $\sigma_{1}=\sigma_{\min}$, we need to choose $H$ such that
    $\sigma_{\min}=\frac{e^{H_{1}}}{e^{H_{0}}}\quad\Longleftrightarrow\quad H_{1}-H_{0}=\log(\sigma_{\min}).$
    By the mean-value-theorem there is a $t^{*}\in[0,1]$ such that $H_{t^{*}}^{\prime}=h_{t^{*}}=\log(\sigma_{\min})$.
  2. 2.
    By change of variables we have that
    $\int_{0}^{1}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t=-\int_{0}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\;\mathrm{d}t=\int_{\sigma_{\min}}^{1}\frac{1}{u}\;\mathrm{d}u=\log(\sigma_{\min}^{-1}).$
Proof of Theorem˜4.5.
For small $t$, we can use the following simple bound, which is independent from the decay of $\operatorname{Cov}(Y^{x,t})$ in $t$:
Lemma 7.2.
Let $t^{*}$ be such that $\sigma_{t^{*}}=\frac{1}{\vartheta}$ for $\vartheta\in\mathbb{R}_{\geq 1}$. Grant ˜4.4 (III). Then
$\int_{0}^{t^{*}}\Gamma_{t}\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta)).$
For large $t$, we need to assume that $\operatorname{Cov}(Y^{x,t})_{ij}$ decays fast enough for $i\neq j$ and $t\rightarrow 1$ to bound the integral over all $B_{ij}$ for $i\neq j$. Under ˜4.4 (I), we know that for all $i\neq j$ and all $x\in\mathbb{R}^{d}$
$\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})$ $\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ij}|\;\mathrm{d}t\leq(t^{*})^{-2\gamma}\int_{t^{*}}^{1}\big{|}(\gamma t^{\gamma-1}\sigma_{t}-\sigma_{t}^{\prime}t^{\gamma})\big{|}\;\mathrm{d}t$
$\leq(t^{*})^{-2\gamma}\gamma\int_{0}^{1}\sigma_{t}-(t^{*})^{-2\gamma}\int_{0}^{1}\sigma_{t}^{\prime}\;\mathrm{d}t=(t^{*})^{-2\gamma}(\gamma+1-\sigma_{\min}).$
For the bound of
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Var}(Y^{x,t}_{i})|\;\mathrm{d}t$
we need to use ˜4.4 (II). Inserting the expression for the variance $\operatorname{Var}(Y^{x,t}_{i})$ we obtain
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Var}(Y^{x,t}_{i})|\;\mathrm{d}t$ $=\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}\big{(}1+O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)})|\;\mathrm{d}t$
$\leq\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}|\;\mathrm{d}t$
$\quad+\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}|\;\mathrm{d}t.$
The first term simplifies to
$\int_{t^{*}}^{1}|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}|\;\mathrm{d}t$ $=\int_{t^{*}}^{1}|\gamma t^{-1}|\;\mathrm{d}t=-\gamma\log(t^{*}).$
For the second term we obtain
$\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{2}O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}|\;\mathrm{d}t$ $\leq\int_{t^{*}}^{1}|(\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}})||\frac{1}{t^{\gamma}}O\big{(}\big{(}\frac{\sigma_{t}}{t^{\gamma}}\big{)}^{\frac{1}{\kappa}}\big{)}|\;\mathrm{d}t$
$\lesssim\int_{t^{*}}^{1}\gamma t^{-1-\frac{\kappa}{\gamma}}\sigma_{t}^{\frac{1}{\kappa}}\;\mathrm{d}t+\int_{t^{*}}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\frac{\sigma_{t}^{\frac{1}{\kappa}}}{t^{\frac{\gamma}{\kappa}}}\;\mathrm{d}t$
$\leq\gamma\int_{t^{*}}^{1}t^{-1-\frac{\kappa}{\gamma}}\;\mathrm{d}t+(t^{*})^{-\frac{\gamma}{\kappa}}\int_{t^{*}}^{1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\sigma_{t}^{\frac{1}{\kappa}}\;\mathrm{d}t$
$=\frac{\gamma^{2}}{\kappa}((t^{*})^{-\frac{\kappa}{\gamma}}-1)+(t^{*})^{-\frac{\gamma}{\kappa}}\kappa(\sigma_{\min}^{\frac{1}{\kappa}}-\sigma_{t^{*}}^{\frac{1}{\kappa}}).$
Using that $\sigma_{\min}^{\frac{1}{\kappa}}-\sigma_{t^{*}}^{\frac{1}{\kappa}}\leq 1$ yields the following bound on the integral over the Lipschitz constant
$\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta))+(t^{*})^{-2\gamma}(\gamma+1)+\frac{\gamma^{2}}{\kappa}((t^{*})^{-\frac{\kappa}{\gamma}}-1)+(t^{*})^{-\frac{\gamma}{\kappa}}\kappa.$
Proof of Theorem˜4.6.
For the proof we need the following additional definition:
Definition 7.3.
A distribution $\mathbb{P}^{*}$ on $\mathbb{R}^{d}$ with density $p^{*}$ satisfies the Poincaré inequality with Poincaré constant $\rho>0$ if for all smooth functions $f$ such that the terms below are well-defined
$\operatorname{Var}(f(X))\coloneqq\int\left(f(x)-\mathbb{E}[f(X)]\right)^{2}p^{*}(x)\mathrm{d}x\leq\frac{1}{\rho}\mathbb{E}[|\nabla f(X)|^{2}].$
From (Boucheron et al., 2013, Theorem 3.20) we know that $\mathcal{N}(0,I_{d})$ satisfies the Poincaré inequality with Poincaré constant $1.$ Thus, via defining $g(x)\coloneqq f(\sigma x+\mu)$ and
$\operatorname{Var}(g(X))\leq\sigma^{2}\mathbb{E}[|\nabla g(X)|^{2}],$
we see that the Poincaré constant of $\mathcal{N}(\mu,\sigma^{2}I_{d})$ is $\sigma^{-2}$. Now we are ready to proof Theorem˜4.6.
Proof.
Property (III):
For the uniform bound, we use that the distribution with density
$q\propto\exp\big{(}-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}\big{)}=\exp\Big{(}-\frac{1}{2}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}\big{(}|y|^{2}-\langle\frac{t}{\sigma_{t}^{2}}\frac{x}{1+\frac{t^{2}}{\sigma_{t}^{2}}},y\rangle\big{)}\Big{)}$ (20)
is a Gaussian distribution with variance $\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}^{-1}I_{d}$. Hence the density defined above satisfies the Poincaré inequality with constant $\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}$. Using the Holley-Strooke perturbation principle Holley and Stroock (1987), in the form of (Ledoux, 2001, Lemma 1.2), we can bound the Poincaré constant $\rho$ of the perturbed Gaussian via
$\rho\geq e^{-4L}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}.$
Thus
$\operatorname{Var}(Y^{x,t}_{i})\leq\big{(}e^{-4L}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}\big{)}^{-1}\leq\frac{e^{4L}}{1+\frac{t^{2}}{\sigma_{t}^{2}}}\leq e^{4L},$
with $L$ from (14). Using
$\operatorname{Cov}(Y^{x,t}_{i},Y^{x,t}_{j})\leq\sqrt{\operatorname{Var}(Y^{x,t}_{i})\operatorname{Var}(Y^{x,t}_{j})},$
we conclude that for all $t\in[0,1]$
$\operatorname{Cov}(Y^{x,t}_{i},(Y^{x,t}_{j})\leq e^{4L}.$
Hence we can set $C=e^{4L}$.
Property (II):
For the variance, we use that
$\operatorname{Var}(Y^{x,t}_{i})=\mathbb{E}[(Y^{x,t}_{i})^{2}]-\mathbb{E}[Y^{x,t}_{i}]^{2}.$
Let $\varphi$ denote the density of $\mathcal{N}(0,I_{d})$. Further, let $t>0$. Then
$\mathbb{E}[Y^{x,t}_{i}]$ $=\frac{\int y_{i}\varphi\big{(}\frac{x-ty}{\sigma_{t}}\big{)}p^{*}(y)\;\mathrm{d}y}{\int\varphi\big{(}\frac{x-ty}{\sigma_{t}}\big{)}p^{*}(y)\;\mathrm{d}y}=\frac{\int\big{(}\frac{x_{i}-z_{i}\sigma_{t}}{t}\big{)}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}=\frac{x_{i}}{t}-\frac{\sigma_{t}}{t}\frac{\int z_{i}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}.$
Similarly
$\mathbb{E}[\big{(}Y^{x,t}_{i}\big{)}^{2}]$ $=\frac{\int\big{(}\frac{x_{i}-z_{i}\sigma_{t}}{t}\big{)}^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}$
$=\big{(}\frac{x_{i}}{t}\big{)}^{2}-\frac{2x_{i}\sigma_{t}}{t^{2}}\frac{\int z_{i}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}+\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\frac{\int z_{i}^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}y}.$
Now define
$A(z_{i})\coloneqq\frac{\int z_{i}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z},\quad A(z_{i}^{2})\coloneqq\frac{\int z_{i}^{2}\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z}.$
We obtain for the variances
$\mathbb{E}[(Y^{x,t}_{i})^{2}]-\mathbb{E}[Y^{x,t}_{i}]^{2}$ $=\big{(}\frac{x_{i}}{t}\big{)}^{2}-\frac{2x_{i}\sigma_{t}}{t^{2}}A(z_{i})+\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}A(z_{i}^{2})-\big{(}\frac{x_{i}}{t}-\frac{\sigma_{t}}{t}A(z_{i})\big{)}^{2}$
$=\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\big{(}A(z_{i}^{2})-A(z_{i})^{2}\big{)}.$ (21)
Hence we need to bound the component variances of a random variable $Z$ with density
$p_{Z}(z)$ $=\frac{\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})}{\int\varphi(z)p^{*}(\frac{x-z\sigma_{t}}{t})\;\mathrm{d}z}$
$=\frac{\exp\big{(}-\frac{|z|^{2}}{2}-\frac{|x-z\sigma_{t}|^{2}}{2t^{2}}-a(\frac{x-z\sigma_{t}}{t})\big{)}}{\int\exp\big{(}-\frac{|z|^{2}}{2}-\frac{|x-z\sigma_{t}|^{2}}{2t^{2}}-a(\frac{x-z\sigma_{t}}{t})\big{)}\;\mathrm{d}z}$
$=\frac{\exp\big{(}-\frac{1}{2}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}|z|^{2}-2\langle z,x\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\rangle\big{)}-a(\frac{x-z\sigma_{t}}{t})\big{)}}{\int\exp\big{(}-\frac{1}{2}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}|z|^{2}-2\langle z,x\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\rangle\big{)}-a(\frac{x-z\sigma_{t}}{t})\big{)}\;\mathrm{d}z}.$ (22)
First we bound the influence of the perturbation function $a$ on the expected value. To do so, we note that
$\mathbb{E}_{p_{Z}}[\partial_{z_{i}}\log(p(Z))]$ $=\int\partial_{z_{i}}p(z)\;\mathrm{d}z$
$=\int...\int_{z_{i}\in\mathbb{R}}\partial_{z_{i}}p(z)\;\mathrm{d}z_{i}\;\mathrm{d}(z_{1},...,z_{i-1},z_{z+1},...,z_{d})$
$=0.$
This implies for (22) that
$0=-\mathbb{E}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}-\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-Z\sigma_{t}}{t})\Big{]}$
$\Longleftrightarrow\quad$ $\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}=\mathbb{E}\Big{[}\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-Z\sigma_{t}}{t})\Big{]}$
$\Longleftrightarrow\quad$ $\Big{|}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{|}=\Bigg{|}\frac{\frac{\sigma_{t}}{t}\mathbb{E}\Big{[}\partial_{z_{i}}a(\frac{x-Z\sigma_{t}}{t})\Big{]}}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}}\Bigg{|}.$
Hence
$\Big{|}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{|}\leq\frac{\sigma_{t}}{t}\frac{L}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}},$ (23)
where $L$ is from (14). From (22) we can see that this bounds the influence of the perturbation function $a$ on the expected value.
Now we bound the variance. To do so, we first note that
$\mathbb{E}_{p_{Z}}[(\partial_{z_{i}}\log(p(Z)))^{2}]$ $=\int\frac{(\partial_{z_{i}}p_{Z}(z))^{2}}{p_{Z}(z)}\;\mathrm{d}z.$ (24)
As
$\partial_{z_{i}}\log(p_{Z}(z))=\frac{\partial_{z_{i}}p_{Z}(z)}{p_{Z}(z)},\quad\partial^{2}_{z_{i}}\log(p_{Z}(z))=\frac{\partial^{2}_{z_{i}}p_{Z}(z)}{p_{Z}(z)}-\frac{(\partial_{z_{i}}p_{Z}(z))^{2}}{(p_{Z}(z))^{2}},$
we have for (24)
$\mathbb{E}_{p_{Z}}[(\partial_{z_{i}}\log(p(Z)))^{2}]$ $=\mathbb{E}_{p_{Z}}\Big{[}\frac{\partial^{2}_{z_{i}}p_{Z}(z)}{p_{Z}(z)}\Big{]}-\mathbb{E}_{p_{Z}}\big{[}\partial^{2}_{z_{i}}\log(p_{Z}(Z))\big{]}.$
Since $\partial_{z_{i}}p_{Z}(z)\rightarrow 0$ for $|z|\rightarrow\infty$, we conclude
$\mathbb{E}_{p_{Z}}[(\partial_{z_{i}}\log(p(Z)))^{2}]$ $=-\mathbb{E}_{p_{Z}}[\partial^{2}_{z_{i}}\log(p_{Z}(Z))].$ (25)
For the right hand side, we obtain
$-\mathbb{E}_{p_{Z}}[\partial^{2}_{z_{i}}\log(p_{Z}(Z))]$ $=\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}+(\frac{\sigma_{t}}{t})^{2}\mathbb{E}_{p_{Z}}\big{[}\partial^{2}_{z_{i}}a(\frac{x-\sigma_{t}z}{t})\big{]}.$
The left side of (25) can be rewritten as
$\mathbb{E}_{p_{Z}}[(\partial_{z_{i}}\log(p(Z)))^{2}]=\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})+\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{)}^{2}\Big{]}$
$=\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\Big{)}^{2}\Big{]}+2\mathbb{E}_{p_{Z}}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$\quad+\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{)}^{2}\Big{]}.$
Then
$\mathbb{E}_{p_{Z}}$ $\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\Big{)}^{2}\Big{]}=\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-\mathbb{E}[Z_{i}]+\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\Big{)}^{2}\Big{]}$
$=\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}(Z_{i}-\mathbb{E}[Z_{i}])\Big{)}^{2}\Big{]}+2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\mathbb{E}_{p_{Z}}\Big{[}(Z_{i}-\mathbb{E}[Z_{i}])\Big{]}$
$\quad+\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{)}^{2}$
$=\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}(Z_{i}-\mathbb{E}[Z_{i}])\Big{)}^{2}\Big{]}+\Big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{)}^{2}\Big{)}.$
Additionally
$\mathbb{E}_{p_{Z}}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}(Z_{i}-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$=\mathbb{E}_{p_{Z}}\Big{[}\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\big{(}Z_{i}-\mathbb{E}[Z_{i}]+\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}\frac{\sigma_{t}}{t}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$=\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}\mathbb{E}_{p_{Z}}\Big{[}(Z_{i}-\mathbb{E}[Z_{i}])\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$\quad+\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}\big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\big{)}\mathbb{E}_{p_{Z}}\Big{[}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}.$
Hence
$\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}Z_{i}-\mathbb{E}[Z_{i}]\Big{)}^{2}\Big{]}$ $=\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}+(\frac{\sigma_{t}}{t})^{2}\mathbb{E}_{p_{Z}}[\partial^{2}_{z_{i}}a(\frac{x-\sigma_{t}z}{t})]$
$-\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{)}^{2}\Big{]}-\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})}\Big{)}^{2}$
$\quad-2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}\mathbb{E}_{p_{Z}}\Big{[}(Z_{i}-\mathbb{E}[Z_{i}])\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}$
$\quad-2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}(\mathbb{E}[Z_{i}]-x_{i}\frac{\sigma_{t}}{t^{2}(1+(\frac{\sigma_{t}}{t})^{2})})\mathbb{E}_{p_{Z}}\Big{[}\partial_{z_{i}}a(\frac{x-\sigma_{t}Z}{t})\Big{]}.$
Combining this with (23) and the fact that $a$ has bounded derivatives, we obtain the following upper and lower bounds:
$\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}Z_{i}-\mathbb{E}[Z_{i}]\Big{)}^{2}\Big{]}$ $\leq\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}+(\frac{\sigma_{t}}{t})^{2}L+2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}L\sqrt{\mathbb{E}_{p_{Z}}\big{[}(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big{]}}$
$\quad+2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}^{2}}{t^{2}}\frac{L^{2}}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}},$
and
$\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\mathbb{E}_{p_{Z}}\Big{[}\Big{(}Z_{i}-\mathbb{E}[Z_{i}]\Big{)}^{2}\Big{]}$ $\geq\big{(}1+(\frac{\sigma_{t}}{t})^{2}\big{)}-(\frac{\sigma_{t}}{t})^{2}L-\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}L^{2}-\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}\Big{(}\frac{\sigma_{t}}{t}\frac{L}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}}\Big{)}^{2}$ (26)
$\quad-2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}}{t}L\sqrt{\mathbb{E}_{p_{Z}}\big{[}(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big{]}}-2\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}\frac{\sigma_{t}^{2}}{t^{2}}\frac{L^{2}}{\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}}.$
With the same reasoning via the Poincaré constant of a Gaussian and the Holley-Stroock perturbation principle, we conclude that
$\mathbb{E}_{p_{Z}}\big{[}(Z_{i}-\mathbb{E}[Z_{i}])^{2}\big{]}\leq e^{4L}.$
As $\big{(}1+\frac{\sigma_{t}^{2}}{t^{2}}\big{)}^{2}>0$ and
$\frac{1}{1+\frac{\sigma_{t}^{2}}{t^{2}}}=1-\frac{\frac{\sigma_{t}^{2}}{t^{2}}}{1+\frac{\sigma_{t}^{2}}{t^{2}}},$
dividing the first term in the upper and lower bound loosens the bound further. Thus we obtain
$\mathbb{E}_{p_{Z}}\big{[}\big{(}Z_{i}-\mathbb{E}[Z_{i}]\big{)}^{2}\big{]}=1+O(\frac{\sigma_{t}}{t}).$
Multiplying with $\frac{\sigma_{t}^{2}}{t}$ yields the result for the variances via (21):
$\operatorname{Var}(Y^{x,t}_{i})=\big{(}\frac{\sigma_{t}}{t}\big{)}^{2}\big{(}1+O\big{(}\frac{\sigma_{t}}{t}\big{)}).$
Property (I):
For the covariances, we need to find a stricter upper bound. We are going to use (Menz, 2014, Theorem 2.3) applied to the component functions $f_{i}\colon\mathbb{R}^{d}\rightarrow\mathbb{R},\;f_{i}(y)=y_{i}$. First we verify that the assumptions are fulfilled in our setting. By construction, the distribution of $Y_{i}^{x,t}$ is a bounded perturbation of a Gaussian. Hence Assumption 2.2 in Menz (2014) is satisfied.
For the Poincaré constant of the $i$-th conditional measure, e.g. the Poincaré constant of the distribution with the density
$p(y_{i}|y_{1},...,y_{i-1},y_{i+1},...,y_{d})\propto\exp\Big{(}-\frac{(x_{i}-ty_{i})^{2}}{2\sigma_{t}^{2}}-\frac{y_{i}^{2}}{2}-a(y_{i}|y_{1},...,y_{i-1},y_{i+1},...,y_{d})\Big{)},$
where $a(\cdot|y_{1},...,y_{i-1},y_{i+1},...,y_{d})\coloneqq a(y_{1},...,y_{i-1},\cdot,y_{i+1},...,y_{d})$ and $i\in\{1,...,d\}$, we can use the same arguments as in the proof of property (III) to obtain
$\rho_{i}^{t}\geq e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}}).$
Note that $a(\cdot|y_{1},...,y_{i-1},y_{i+1},...,y_{d})$ is still bounded. Furthermore, we know that the off-diagonal entries of the Hessian of the $\log$-density are bounded by $L$. To profit from easier notation later, we define the matrix
$A_{t}=(A_{t,ij})_{i,j=1,...,d},\quad\text{where}\quad A_{t,ij}\coloneqq\begin{cases}\rho^{t}_{i}-e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}},&i=j,\\ -L,&i\neq j.\end{cases}$
We need to find a bound on $t^{*}$ such that the matrix $A_{t}+e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}$ is positive definite. As $A_{t}$ is symmetric by construction, we know that the eigenvalues of $A_{t}+e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}$ are real numbers. By Gerschgorins theorem (Gerschgorin, 1931, Satz 2), we know that the eigenvalues of $A_{t}+e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}$ are in the following union of intervals:
$D=\bigcup_{i=1}^{d}\Big{[}\rho_{i}^{t}-L(d-1),\rho_{i}^{t}+L(d-1)\Big{]}.$ (27)
Inserting the lower bound of $\rho_{i}^{t}$, we choose $t^{*}$ such that for all $t<t^{*}$
$e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}})>L(d-1)\quad\Longleftarrow\quad\frac{\sigma_{t}}{t}<\frac{1}{\sqrt{e^{4L}L(d-1)-1}}.$ (28)
For this choice of $t^{*}$, we conclude that the matrix $A_{t,ij}+e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}$ is positive definite. Thus all of the assumptions in (Menz, 2014, Theorem 2.3) are satisfied.
We conclude that
$|\operatorname{Cov}(Y^{x,t})_{ij}|\leq((e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}+A_{t})^{-1})_{ij}.$
Now
$\Big{(}e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}I_{d}+A_{t}\Big{)}^{-1}$ $=e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}\Big{(}I_{d}+e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t}\Big{)}^{-1}.$ (29)
As $A_{t}$ is a symmetric matrix, the spectral norm is the absolute value of the largest eigenvalue. Using Gerschgorins theorem (Gerschgorin, 1931, Satz II) again, we know that the eigenvalues of $A_{t}$ are in the following union of intervals:
$D=\bigcup_{i=1}^{d}\Big{[}\rho_{i}^{t}-e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}-L(d-1),\rho_{i}^{t}-e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}+L(d-1)\Big{]}.$ (30)
As $\rho_{i}^{t}-e^{-4L}\frac{t^{2}}{\sigma_{t}^{2}}=e^{-4L}$ we conclude that the largest eigenvalue is bounded by
$\lambda_{\max}(A_{t})\leq e^{-4L}+L(d+1).$
We therefore deduce that
$\Big{\|}e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t}\Big{\|}=e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}\Big{\|}A_{t}\Big{\|}\leq\frac{\sigma_{t}^{2}}{t^{2}}(1+e^{4L}L(d-1)).$
If we choose $t^{*}$ such that for $t>t^{*}$
$\frac{\sigma_{t}^{2}}{t^{2}}(1+e^{4L}L(d-1))\leq\frac{\sigma_{t}}{t}<1\quad\Longleftrightarrow\quad\frac{\sigma_{t}}{t}\leq\frac{1}{1+e^{4L}L(d-1)},$
we can use a Neumann series. This gives
$\Big{(}I_{d}+e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t}\Big{)}^{-1}=\sum_{k=0}^{\infty}(-1)^{k}e^{4Lk}\frac{\sigma_{t}^{2k}}{t^{2k}}A_{t}^{k}=I_{d}-e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t}+\sum_{k=2}^{\infty}(-1)^{k}e^{4Lk}\frac{\sigma_{t}^{2k}}{t^{2k}}A_{t}^{k}.$
Now for the $ij$-th element, we obtain
$\Big{(}I_{d}+e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t}\Big{)}^{-1}_{ij}\leq\Big{(}\mathds{1}_{i=j}+e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t,ij}+\Big{\|}\sum_{k=2}^{\infty}(-1)^{k}e^{4Lk}\frac{\sigma_{t}^{2k}}{t^{2k}}A_{t}^{k}\Big{\|}\Big{)}.$
Then we can bound
$\Big{\|}\sum_{k=2}^{\infty}(-1)^{k}e^{4Lk}\frac{\sigma_{t}^{2k}}{t^{2k}}A_{t}^{k}\Big{\|}$ $\leq\sum_{k=2}^{\infty}\Big{\|}e^{4Lk}\frac{\sigma_{t}^{2k}}{t^{2k}}A_{t}^{k}\Big{\|}\leq\sum_{k=2}^{\infty}\Big{\|}e^{4L}\frac{\sigma_{t}^{2}}{t^{2}}A_{t}\Big{\|}^{k}\leq\sum_{k=2}^{\infty}\big{(}\frac{\sigma_{t}}{t}\big{)}^{k}$
Using the convergence of the geometric series, we get that
$\sum_{k=2}^{\infty}\big{(}\frac{\sigma_{t}}{t}\big{)}^{k}=\frac{\frac{\sigma_{t}^{2}}{t^{2}}}{1-\frac{\sigma_{t}}{t}}\leq\frac{\sigma_{t}^{2}}{t^{2}}.$
Inserting everything into (29), we obtain
$\Big{(}I_{d}+\frac{\sigma_{t}^{2}}{t^{2}}A_{t}\Big{)}^{-1}_{ij}\lesssim\frac{\sigma_{t}^{2}}{t^{2}}(\mathds{1}_{i=j}+\frac{\sigma_{t}^{2}}{t^{2}}).$
If we choose $t^{*}\geq\frac{1}{2}$, we know that $\frac{\sigma_{t}^{2}}{t^{2}}\leq 2\sigma_{t}^{2}$. This gives the smaller bound on the covariances.
7.3 Proofs of Section˜5
Proof of Theorem˜5.2.
To validate the bound on $|v_{t}^{j}|_{\infty}$, we insert $\log(\sigma_{\min})\sim\log(n)$ a bit later. We start by defining $A\coloneqq[-\log(n),\log(n)]^{d}$ and
$g\colon\mathcal{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},\quad g(v,y)\coloneqq\int_{0}^{1}\int|v_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t.$ (31)
Then we use Lemma˜2.1,
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\hat{v}_{t}(X_{t})|^{2}\big{]}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\hat{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}-C,$ (32)
where
$C\coloneqq\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}[|v_{t}(X_{t})-v_{t}(X_{t}|Y)|^{2}].$
We can split the integral
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\hat{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}$ $=\int_{A}\int_{0}^{1}\int|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(y)\;\mathrm{d}y$
$\quad+\int_{\mathbb{R}^{d}\setminus A}\int_{0}^{1}\int|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(y)\;\mathrm{d}y.$
Note that
$\int_{A}\int_{0}^{1}\int|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(x)\;\mathrm{d}y=\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\hat{v},Y)\mathds{1}_{Y\in A}].$
We have for every $x\in\mathbb{R}^{d}$ and $j\in\{1,...,d\}$ and the $t$-th component function $v_{t}^{j}$ of $v_{t}$
$|v_{t}^{j}(x)|$ $=\Big{|}\int v^{j}_{t}(x|y)\frac{p_{t}(x|y)}{\int p_{t}(x|z)p^{*}(z)\;\mathrm{d}z}p^{*}(y)\;\mathrm{d}y\Big{|}$
$=\Big{|}\int\big{(}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}x_{j}+(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t)y_{j}\big{)}\frac{p_{t}(x|y)}{\int p_{t}(x|z)p^{*}(z)\;\mathrm{d}z}p^{*}(y)\;\mathrm{d}y\Big{|}$
$\leq\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}|x_{j}|+\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}\frac{\int|y_{j}|\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}$
$\leq\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}|x_{j}|+e^{2L}\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}\mathbb{E}_{Z\sim\mathcal{N}(\frac{t}{t^{2}+\sigma_{t}^{2}}x,(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-1}I_{d})}[|Z_{j}|]$
$\leq\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}|x_{j}|+e^{2L}\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}(\frac{t}{t^{2}+\sigma_{t}^{2}}|x_{j}|+(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}})$
$=\big{(}\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}+e^{2L}\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}\frac{t}{t^{2}+\sigma_{t}^{2}}\big{)}|x_{j}|+\big{|}\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}\big{|}(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}}.$
As
$\max_{t\in[0,1]}\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}=\log(\sigma_{\min}^{-1}),\quad(1+\frac{t^{2}}{\sigma_{t}^{2}})^{-\frac{1}{2}}\leq 1$
and due to Lemma˜7.10
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\max(\log(\sigma_{\min}^{-1}),e^{2})$
we obtain for the maximum over $t\in[0,1]$ for $n$ big enough
$|v_{t}^{j}(x)|$ $\lesssim\log(n)^{2}|x_{j}|+\log(n).$
The constants lead to the bound in (15). An analogous calculation shows that, for $t\in[0,1]$, the norm of $v_{t}$ is bounded by
$|v_{t}(x)|$ $\lesssim\log(n)^{2}|x|+\log(n).$ (33)
Therefore, we obtain for every $v\in\mathcal{M}$ and $y\in A$ using the construction of $v_{t}(\cdot|\cdot)$ and fact that the functions in $\mathcal{M}$ are cut at (15) $\log(n),\log(n)]^{d}$ for all $t$,
$g(v,y)$ $=\int_{0}^{1}\int|v_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t$
$\lesssim\log(n)^{6}+\int_{0}^{1}\int|v(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t.$
Since
$\int|v(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x$ $=\int\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}x+\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}y\big{|}^{2}p_{t}(x|y)\;\mathrm{d}x$
$=\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}^{2}\int|x|^{2}p_{t}(x|y)\;\mathrm{d}x+\big{(}1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t\big{)}^{2}|y|^{2},$
and as $p_{t}(\cdot|y)$ is the density of a Gaussian with mean $y$ and variance $\sigma_{t}^{2}I_{d}$,
$\int|x|^{2}p_{t}(x|y)\;\mathrm{d}x$ $=t^{2}|y|^{2}+\sigma_{t}^{2}d,$ (34)
we obtain for $g(v,y)$
$g(v,y)$ $\lesssim\log(n)^{6}+\int_{0}^{1}\big{|}\frac{\sigma^{\prime}_{t}}{\sigma_{t}}\big{|}^{2}(t^{2}|y|^{2}+\sigma_{t}^{2}d)\;\mathrm{d}t+|y|^{2}\int_{0}^{1}(1-\frac{\sigma^{\prime}_{t}}{\sigma_{t}}t)^{2}\;\mathrm{d}t\lesssim\log(n)^{6}.$
We denote the constant in the bound of $g(v,y)$ by $D$. Now we can use the Bernstein-type concentration inequality from Chen et al. (2023b, Lemma 15) and conclude that for every $a\in(0,1],\delta_{1}\in(0,\frac{1}{3})$ and $\tau\in\mathbb{R}_{>0}$
$\mathbb{P}\Big{(}\sup_{\bar{v}\in\mathcal{M}}$ $\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\bar{v},Y)\mathds{1}_{Y\in A}]-\frac{1+a}{n}\sum_{i=1}^{n}g(\bar{v},X_{i})\mathds{1}_{X_{i}\in A}$
$\quad>\frac{(1+6/a)D\log(n)^{6}}{3n}\log\Big{(}\frac{\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})}{\delta_{1}}\Big{)}+(1+a)\tau\Big{)}\leq\delta_{1}.$
We keep the term $\frac{a}{n}\sum_{i=1}^{n}g(\bar{v},X_{i})\mathds{1}_{X_{i}\in A}$ separate. For $\frac{1}{n}\sum_{i=1}^{n}g(\bar{v},X_{i})\mathds{1}_{X_{i}\in A}$ we conclude that with a probability of at least $1-\delta_{1}$
$\int_{A}\int\int$ $|\hat{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(x)\;\mathrm{d}y$
$\leq\frac{1+a}{n}\sum_{i=1}^{n}g(\hat{v},X_{i})\mathds{1}_{X\in A}+\frac{(1+6/a)D\log(n)^{6}}{3n}\log\Big{(}\frac{\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})}{\delta_{1}}\Big{)}+(1+a)\tau.$
Due to the choice of $\hat{v}$ as the empirical risk minimizer, we know that for every $\tilde{v}\in\mathcal{M}$
$\frac{1}{n}\sum_{i=1}^{n}g(\hat{v},X_{i})$ $=\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\in A}$
$=\frac{1}{n}\sum_{i=1}^{n}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\notin A}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\int\int|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t$
$\quad-\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\hat{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\notin A}.$
As
$\frac{1}{n}\sum_{i=1}^{n}\int\int|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t$ $=\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\in A}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\notin A},$
we can use the other case of Chen et al. (2023b, Lemma 15) to conclude that for every $a\in(0,1],\delta_{2}\in(0,\frac{1}{3})$ and $\tau\in\mathbb{R}_{>0}$
$\mathbb{P}\Big{(}\sup_{\bar{v}\in\mathcal{M}}\frac{1}{n}$ $\sum_{i=1}^{n}g(\bar{v},X_{i})\mathds{1}_{X_{i}\in A}-(1+a)\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\bar{v},Y)\mathds{1}_{Y\in A}]$
$\quad>\frac{(1+3/a)D\log(n)^{6}}{3n}\log\Big{(}\frac{\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})}{\delta_{2}}\Big{)}+(1+a)\tau\Big{)}\leq\delta_{2}.$
We conclude that with a probability of $1-\delta_{2}$
$\frac{1}{n}\sum_{i=1}^{n}\Big{(}\int\int$ $|\tilde{v}_{t}(x)-v_{t}(x|X_{i})|^{2}p_{t}(x|X_{i})\;\mathrm{d}x\;\mathrm{d}t\Big{)}\mathds{1}_{X_{i}\in A}$
$\leq(1+a)\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]+\frac{(1+3/a)D\log(n)^{6}}{3n}\log\Big{(}\frac{\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})}{\delta_{2}}\Big{)}$
$\quad+(1+a)\tau.$
Like before, we separate $a\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]$. Using Lemma˜2.1 again, we obtain
$\mathbb{E}_{Y\sim\mathbb{P}^{*}}[g(\tilde{v},Y)\mathds{1}_{Y\in A}]$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}$
$\quad-\int_{\mathbb{R}^{d}\setminus A}\int\int|\tilde{v}_{t}(x)-v_{t}(x|y)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\;p^{*}(x)\;\mathrm{d}y.$
Further
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ Y\sim p^{*},\\ X_{t}\sim p_{t}\left(\cdot|Y\right)\end{subarray}}\big{[}\left|\tilde{v}_{t}(X_{t})-v_{t}\left(X_{t}|Y\right)\right|^{2}\big{]}=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}+C.$
Note that $C$ cancels with the same constant in (32). The terms arising from the restriction of the Bernstein-bound to the set $A$ can be bounded with high probability by the following result.
Lemma 7.4.
Fix $C\in\mathbb{R}$. Let $\log(\sigma_{\min}^{-1})\sim\log(n)$. With a probability of $1-\frac{1}{3n}$ for $n\geq e^{3(d+7+\frac{1}{2})}$
$|\mathbb{E}[g(\hat{v},Y)\mathds{1}_{Y\notin A}]-\mathbb{E}[g(\tilde{v},Y)\mathds{1}_{Y\notin A}]+\sum_{i=1}^{n}g(\tilde{v},X_{i})\mathds{1}_{X_{i}\notin A}-g(\hat{v},X_{i})\mathds{1}_{X_{i}\notin A}|\leq n^{-\frac{1}{2}}.$
Collecting all other terms, inserting $\log(\sigma_{\min})\sim\log(n)$, setting $\delta_{1}=\delta_{2}=\frac{1}{3n}$ and using the union bound leads to the result. The infimum can be used due to the presence of other nonzero terms. ∎
Preparation for the proof of Theorem˜5.3.
The proof of Theorem˜5.3 requires some additional results to bound the covering number and apply the approximation result. We further need to look carefully into the approximation result by Gühring et al. (2020).
As we cannot approximate a function on $\mathbb{R}^{d}$ with a finite neural network with fixed precision, we are going to look at functions that approximate $v$ on the set $\log(n),\log(n)]^{d}\times[0,1]$ and show that the error on the complement is small. Hence we want to determine a network necessary to obtain
$\int\int_{[-\log(n),\log(n)]^{d}}|\tilde{v}_{t}(x)-v_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t\leq\varepsilon,$
for a given approximation error $\varepsilon>0$. We map any point outside of $\log(n),\log(n)]^{d}$ back to this hypercube using the one layer ReLU net associated with the following function
$\operatorname{clip}_{\text{input}}(x_{i},-\log(n),\log(n))=x_{i}-\phi(x_{i}-\log(n))+\phi(-x_{i}-\log(n)).$
Thus, the Lipschitz constant of a network that approximates $v$ on the set $\log(n),\log(n)]^{d}\times[0,1]$ is bounded by the Lipschitz constant of the network on $\log(n),\log(n)]^{d}\times[0,1]$. A similar one layer clipping function can be used to clip the component functions at (15). The order of the number of layers and the number of nonzero weights will not change with both adaptions. For the error on the complement, we use the following result:
Lemma 7.5.
For every $\tilde{v}$ in $\mathcal{M}$ and $\log(\sigma_{\min}^{-1})\sim\log(n)$, if $n\geq e^{4(\frac{1}{2}+6+d)}$ then
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t\leq n^{-\frac{1}{2}}.$
Bound of the covering number:
Lemma 7.6.
We have for $\bar{v}^{1},\bar{v}^{2}\in\mathcal{M}$ and every $y\in[-\log(n),\log(n)]^{d}$, that
$|g(\bar{v}^{1},y)-g(\bar{v}^{2},y)|\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}\log(n)^{4}.$
Hence we can bound
$\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty})\leq\mathcal{N}(\frac{\tau}{\log(n)^{4}},\mathcal{M},\|\cdot\|_{\infty}).$
From Suzuki (2019, Lemma 3) we know that the covering number of the set ReLU networks of a $0,1]^{d}$, denoted by $\mathsf{NN}_{\text{ReLU}}$ cube is
$\log(\mathcal{N}(\tau,\mathsf{NN}_{\text{ReLU}},\|\cdot\|_{\infty,[0,1]^{d}}))\leq 2SL\log\left(\tau^{-1}L(B\vee 1)(W+1)\right),$
where $L$ is the number of layers, $S$ is the number of nonzero weights, $B$ is the maximal absolute value of a single weight and $W$ is the maximal width. A careful inspection of the proof of the approximation result by Gühring et al. (2020) will later reveal the specific choices.
To consider functions on $\log(n),\log(n)]^{d}\times[0,1]$, we proceed analogously to (Yakovlev and Puchkin, 2025, p.46), multiplying the first $d$ coordinates of the weight matrix of the first lemma with $\log(n)$ and dividing the input vector with the same value. This has of course an impact on the bound of the weights, which will scale up by the factor $\log(n)$. Additionally, the weights of the clipping layers are of order $\log(n)^{3}$. We thus obtain
$\log(\mathcal{N}(\tau,g(\mathcal{M}),\|\cdot\|_{\infty}))\lesssim 2SL\log\left(\tau^{-1}\log(n)^{4}L(B\vee 1)\log(n)^{3}(W+1)\right).$
Approximation result:
From Gühring et al. (2020, Corollary 4.2) we know that for any $f\in C^{s}((0,1)^{d}),s\in\mathbb{N}\setminus\{1\}$ with bounded $s$-th order partial derivatives we can approximate $f$ and its Lipschitz constant simultaneously using a ReLU network as defined in (17). More precisely, for every $\varepsilon\in(0,1/2)$ there exists a ReLU network $f_{\mathsf{NN}}$ such that
$\left\|f_{\mathsf{NN}}-f\right\|_{\mathcal{H}^{1}\left((0,1)^{d}\right)}\leq\varepsilon.$ (35)
The network $f_{\mathsf{NN}}$ can be chosen with at most
$L\leq c\cdot\log\left(\varepsilon^{-\frac{\xi}{\beta-1}}\right)$
hidden layers and
$S\leq c\cdot\varepsilon^{-\frac{d}{s-1}}\cdot\log^{2}\left(\varepsilon^{-\frac{s}{s-1}}\right)$
nonzero weights, where $c$ is a constant.
We remark that Gühring et al. (2020) use the term "standard neural networks" for these ReLU architectures, in contrast to networks with skip connections that they also study. Additionally, note that we need to use $d$ parallel networks in order to approximate the function $v$.
As the upper bound depends on the Lipschitz constant of the network, this stronger approximation result is necessary. Since the covering number depends on the maximum bound of the weights, we need to track this in the proof of Gühring et al. (2020, Theorem 4.1). Gühring et al. (2020, Lemma C.3 (v)) reveals that the biggest single weight is set to $N$, which is later chosen as $\Big{\lceil}\Big{(}\frac{\varepsilon}{2CL}\Big{)}^{-\frac{1}{s-1}}\Big{\rceil}$, where $L$ is the bound on the partial derivatives up to the highest order considered. To apply the result of Gühring et al. (2020, Corollary 4.2) without tracking the bound on the partial derivatives in their proof, we approximate $v$ divided by this bound and scale the approximated function up in the end. This can be achieved adding a layer in front and a layer after the network. The maximal weight used is of the size of the bound.
As $v$ is in $C^{\infty}$, we can apply (Gühring et al., 2020, Corollary 4.2) for arbitrary large $s$. However, the use of a larger $s$ will lead to a larger bound on the absolute value of the partial derivatives up to the order $s$. Since this scales up the approximation error, we need a precise quantification. The next result uses the $\log$-Sobolev inequality to bound higher derivatives of $v$.
Lemma 7.7.
Fix $s\in\mathbb{N}$. Let $\{i_{1},...,i_{s}\}\subset\{1,...,d\}^{s}$. For the $k$-th order partial derivative of $v$ we have that
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)\lesssim\log(\sigma_{\min}^{-1})\sigma_{\min}^{-s+1}.$
For the derivatives with respect to $t$, we obtain
Lemma 7.8.
Fix $s\in\mathbb{N}$. Assume that $\log(\sigma_{\min})\geq s$. For every $k\in\{1,...,s\}$ we have that
$\frac{\partial^{k}}{\partial t^{k}}v^{j}_{t}(x)\lesssim\operatorname{polylog}(\sigma_{\min}^{-1})\operatorname{polylog}(n)\sigma_{\min}^{-s-2}.$
The bound on mixed derivatives follows exactly the same lines as the proofs of Lemma˜7.7 and Lemma˜7.8
Now we are ready to proof Theorem˜5.3. Proof of Theorem˜5.3.
We scale the function down by $\sigma_{\min}^{-s-2}\log^{2}(n)$ and approximate $v\cdot\sigma_{\min}^{-s-2}\log^{2}(n)$ instead of $v$. After that, we add another layer to scale the output of the neural net up. The approximation error $\varepsilon$ will also scale up by $\sigma_{\min}^{-s-2}\log^{2}(n)$. The maximal weight in the inner network will then no longer depend on this bound, but the largest weight might increase. Hence we can set
$W=C(\varepsilon^{-\frac{1}{s-2}}\vee\sigma_{\min}^{-s-2}\log^{2}(n)).$
Inserting everything into Theorem˜5.2 and setting $\tilde{d}=d+1$, we obtain for $a<1$ with a probability of $1-\frac{1}{n}$ for $n$ big enough
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}$
$\lesssim+\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t+(2+a)\tau+a\log(n)^{4}+\varepsilon\sigma_{\min}^{-s-2}\log^{2}(n)$
$\quad+\frac{a^{-1}\log(n)^{7}}{n}\varepsilon^{-\frac{\tilde{d}}{s-1}}\log\big{(}2\big{(}\tau^{-1}\log(n)^{7}L(B\vee 1)(\varepsilon^{-\frac{1}{s-1}}\vee\sigma_{\min}^{-s-2}\log^{2}(n)\big{)}\big{)}$
$\quad+(2+a)\tau+a\log(n)^{6}+\varepsilon\sigma_{\min}^{-s-2}\log^{2}(n).$ (36)
Note that $B$ and $L$ will depend on $\varepsilon$, thus we are restricting feasible choices of $\varepsilon$ to choices that grow at most polynomial in $n^{-1}$. Ignoring logarithmic terms, we first solve for $a$:
$\frac{a^{-1}\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\sim a\quad\Longleftrightarrow\quad a=\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}.$
$\tau$ can be set such that the corresponding term is of the same order, it suffices to set
$\tau\sim\frac{\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}}{2+\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}}.$
This choice, as well as the choice of $\varepsilon$, influence the first term in (36) only logarithmically. Now we solve for $\varepsilon$
$\Big{(}\frac{\varepsilon^{-\frac{\tilde{d}}{s-1}}}{n}\Big{)}^{\frac{1}{2}}\sim\varepsilon\sigma_{\min}^{-s-2}\quad\Longleftrightarrow\quad\varepsilon\sim n^{-\frac{s-1}{\tilde{d}+2s-2}}\sigma_{\min}^{(s+2)\frac{2s-2}{\tilde{d}+2s-2}}.$
Restricting $\sigma_{\min}$ to polynomials in $n$, we ensure $\varepsilon$ is polynomial in $n$. Thus, ignoring logarithmic factors in $n$, we obtain with a probability of $1-\frac{1}{n}$ for $n$ big enough
$\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1]\\ X_{t}\sim p_{t}\end{subarray}}\big{[}|v_{t}(X_{t})-\tilde{v}_{t}(X_{t})|^{2}\big{]}\lesssim\operatorname{polylog}(n)n^{-\frac{s-1}{\tilde{d}+2s-2}}\sigma_{\min}^{(s+2)\big{(}\frac{2s-2}{\tilde{d}+2s-2}-1\big{)}}.$
To choose $\sigma_{\min}$ such that is is optimal in the setting of Theorem˜3.1, we have to set it such that
$\sigma_{\min}^{1+\alpha}\sim\Big{(}n^{-\frac{s-1}{d+2s-2}}\sigma_{\min}^{(s+2)\big{(}\frac{2s-2}{\tilde{d}+2s-2}-1\big{)}}\Big{)}^{\frac{1}{2}}\quad\Longleftrightarrow\quad\sigma_{\min}\sim n^{-\frac{1}{\frac{s+2}{s-1}\tilde{d}+(2\alpha+2)(\frac{\tilde{d}}{s-1}+2)}}.$
This choice is a polynomial of $n^{-1}$ and hence a feasible choice for $\sigma_{\min}$. For $n\geq 1$ and $\sigma_{\min}<1$, we know that the network approximates the Lipschitz constant of $v$ with precision $\varepsilon\sigma_{\min}^{-s-2}$. Plugging in the choice of $\sigma_{\min}$ leads to
$\varepsilon\sigma_{\min}^{-s-2}\sim n^{-\frac{2+2\alpha}{\frac{s+2}{s-1}\tilde{d}+(2\alpha+2)(\frac{\tilde{d}}{s-1}+2)}}\leq 1.$
Hence we obtain for the Lipschitz constant $\hat{L}_{t}$ of $\hat{v}_{t}$:
$e^{\int_{0}^{1}\hat{L}_{t}\;\mathrm{d}t}\leq e^{\int_{0}^{1}\Gamma_{t}+1\;\mathrm{d}t}\leq e^{C+1}.$
Note that we cannot let $s\rightarrow\infty$, as this will blow up the constant. However, for every fixed $\eta>0$, we can choose $s$ such that
$s=\big{\lceil}(5+2\alpha)\frac{\tilde{d}}{\eta}+1\big{\rceil}.$
This leads to the final choice
$\sigma_{\min}\sim n^{-\frac{1}{\tilde{d}+4\alpha+4+\eta}}$
Now we use the results of Lemma˜7.4, set $\delta=\frac{1}{2n}$ and use the union bound. Combining this with Theorem˜3.1, Theorem˜4.5, (Kunkel and Trabs, 2025b, Theorem 3.3) and $\tilde{d}=d+1$, we obtain with a probability of $1-\frac{1}{n}$
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})\lesssim\operatorname{polylog}(n)n^{-\frac{1+\alpha}{d+4\alpha+5+\eta}}.$
The number of hidden layers $L$ of $f_{\mathsf{NN}}$ is bounded by $c\cdot\log\left(n\right)$ and the number of nonzero weights $S$ is bounded by $c\cdot n^{c(d,\alpha,\eta)}\cdot\log^{2}\left(n\right)$, where $c$ is a constant independent from $n$. ∎
7.4 Additional proofs of Section˜7 and helper results
Proof of Lemma˜7.1.
By the definition of $v_{t}$ and $v_{t}(\cdot|\cdot)$, we get
$v_{t}(x)=\int\big{(}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-\mu_{t}(y))+\mu_{t}^{\prime}(y)\Big{)}\frac{p_{t}(x|y)}{p_{t}(x)}p^{*}(y)\;\mathrm{d}y,$
where indicates the derivative with respect to $t.$ For the Jacobian, we get using $\mu_{t}(y)=t^{\gamma}y$
$D_{x}v_{t}(x)=D_{x}\int\big{(}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x-\mu_{t}(y))+\mu_{t}^{\prime}(y)\Big{)}\frac{p_{t}(x|y)}{p_{t}(x)}p^{*}(y)\;\mathrm{d}y$ (37)
$=D_{x}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\frac{x}{p_{t}(x)}\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y-D_{x}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\frac{t^{\gamma}}{p_{t}(x)}\int yp_{t}(x|y)p^{*}(y)\;\mathrm{d}y+D_{x}\frac{\gamma t^{\gamma-1}}{p_{t}(x)}\int yp_{t}(x|y)p^{*}(y)\;\mathrm{d}y$ (38)
$=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}I_{d}+\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}D_{x}\frac{\int yp_{t}(x|y)p^{*}(y)\;\mathrm{d}y}{p_{t}(x)}.$ (39)
For the derivative with respect to $x_{i}$ of the jthj-th coordinate function, we get using the dominated convergence theorem
$\frac{\partial}{\partial x_{i}}\frac{\int y_{j}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}{\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}$
$=\frac{(\int y_{j}\frac{\partial}{\partial x_{i}}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)-(\int y_{j}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int\frac{\partial}{\partial x_{i}}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)}{(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)^{2}}$
$=\frac{(\int y_{j}(-\frac{x_{i}-t^{\gamma}y_{i}}{\sigma_{t}^{2}})p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)-(\int y_{j}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int(-\frac{x_{i}-t^{\gamma}y_{i}}{\sigma_{t}^{2}})p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)}{(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)^{2}}$
$=\frac{t^{\gamma}\int y_{j}y_{i}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}{\sigma_{t}^{2}\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}-\frac{t^{\gamma}(\int y_{j}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)(\int y_{i}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)}{\sigma_{t}^{2}(\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y)^{2}}.$
For fixed $x$ and $t$ let $Y^{x,t}$ be a random variable with density $\frac{p_{t}(x|\cdot)p^{*}(\cdot)}{\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}.$ Then for $t\in[0,1)$
$\frac{\partial}{\partial x_{i}}\frac{\int y_{j}p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}{\int p_{t}(x|y)p^{*}(y)\;\mathrm{d}y}=\frac{t^{\gamma}}{\sigma_{t}^{2}}(\mathbb{E}[Y_{i}^{x,t}Y_{j}^{x,t}]-\mathbb{E}[Y_{i}^{x,t}]\mathbb{E}[Y_{j}^{x,t}])=\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ji}.$
We obtain for the derivative of with respect to $x_{i}$ of the jthj-th coordinate function
$\frac{\partial}{\partial x_{i}}v^{j}_{t}(x)=\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{x,t})_{ji}.$
Proof of Lemma˜7.2.
Let $t^{*}$ be such that $\sigma_{t^{*}}=\frac{1}{\vartheta}$. Then
$\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\mathds{1}_{i=j}+\big{(}\gamma t^{\gamma-1}$ $-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}\;\mathrm{d}t$
$\leq\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t+\int_{0}^{t^{*}}\big{|}\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}\;\mathrm{d}t.$
For the first term, we can use that $\sigma_{t}^{\prime}$ is non positive and hence by change of variables
$\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t=-\int_{0}^{t^{*}}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\;\mathrm{d}t=\int_{\sigma_{t^{*}}}^{1}\frac{1}{u}\;\mathrm{d}u=\log(\sigma_{t^{*}}^{-1})=\log(\vartheta).$
For the second integral, we can bound the Covariance term by
$\big{|}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}$ $\leq C.$
Now we get
$\int_{0}^{t^{*}}$ $\big{|}\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\operatorname{Cov}(Y^{\cdot,t})_{ij}\big{|}\;\mathrm{d}t\leq C\int_{0}^{t^{*}}\big{|}\big{(}\gamma t^{\gamma-1}-\frac{\sigma_{t}^{\prime}t^{\gamma}}{\sigma_{t}}\big{)}\frac{t^{\gamma}}{\sigma_{t}^{2}}\big{|}\;\mathrm{d}t$
$\leq C\int_{0}^{t^{*}}\Big{|}\gamma t^{\gamma-1}\frac{t^{\gamma}}{\sigma_{t}^{2}}\Big{|}\;\mathrm{d}t+C\int_{0}^{t^{*}}\Big{|}\frac{\sigma_{t}^{\prime}t^{2\gamma}}{\sigma^{3}_{t}}\Big{|}\;\mathrm{d}t\leq\vartheta^{2}C\int_{0}^{t^{*}}\gamma t^{\gamma-1}\;\mathrm{d}t+\vartheta^{2}C\int_{0}^{t^{*}}\big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\big{|}\;\mathrm{d}t\lesssim\vartheta^{2}(1+\log(\vartheta)).$
Proof of Lemma˜7.4.
First we abbreviate
$B\coloneqq\mathbb{E}[g(\hat{v},Y)\mathds{1}_{Y\notin A}]-\mathbb{E}[g(\tilde{v},Y)\mathds{1}_{Y\notin A}]+\sum_{i=1}^{n}g(\tilde{v},X_{i})\mathds{1}_{X_{i}\notin A}-g(\hat{v},X_{i})\mathds{1}_{X_{i}\notin A}.$
[…]
$g(v,y)$ $\lesssim\log(n)^{6}+\log(n)^{2}|y|^{2}+\log(\sigma_{\min}^{-1})^{2}\log(n)^{2}|y|^{2}.$
Thus, we obtain
$\mathbb{E}_{X_{i}}[|B|]$ $\lesssim\mathbb{E}\Big{[}\Big{|}\mathds{1}_{Y\notin A}\big{(}\log(n)^{6}+\log(n)^{2}|Y|^{2}\big{)}\Big{|}\Big{]}=\log(n)^{6}\int_{\mathbb{R}^{d}\setminus A}p^{*}(y)\;\mathrm{d}y+\log(n)^{2}\int_{\mathbb{R}^{d}\setminus A}|y|^{2}p^{*}(y)\;\mathrm{d}y.$
Using the structure of $p^{*}$ and assuming $n\geq 4$, we obtain
$\int_{\mathbb{R}^{d}\setminus A}p^{*}(y)\;\mathrm{d}y$ $\leq\int_{\mathbb{R}^{d}\setminus A}|y|^{2}p^{*}(y)\;\mathrm{d}y=\frac{\int_{\mathbb{R}^{d}\setminus A}|y|^{2}\exp(-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\exp(-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}$
$\leq e^{2L}\frac{\int_{\mathbb{R}^{d}\setminus A}|y|^{2}\exp(-\frac{|y|^{2}}{2})\;\mathrm{d}y}{\int\exp(-\frac{|y|^{2}}{2})\;\mathrm{d}y}=e^{2L}\Big{(}2\frac{\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}{\int\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}\Big{)}^{d}.$
As
$0\leq\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y=[-y_{1}\exp(-\frac{y_{1}^{2}}{2})]_{\log(n)}^{\infty}-\int_{\log(n)}^{\infty}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}\leq\log(n)\exp(-\frac{\log(n)^{2}}{2}),$
we can bound the above by
$e^{2L}\Big{(}2\frac{\int_{\log(n)}^{\infty}y_{1}^{2}\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}{\int\exp(-\frac{y_{1}^{2}}{2})\;\mathrm{d}y_{1}}\Big{)}^{d}$ $\lesssim\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}.$
For large $n$, this decays faster than any polynomial in $n$. We conclude
$\mathbb{E}_{X_{i}}[|B|]\lesssim\log(n)^{6}\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}.$
Now Markov’s inequality yields
$\mathbb{P}(B\geq n^{-\frac{1}{2}})$ $\lesssim\log(n)^{6}\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}n^{\frac{1}{2}},$
$\mathbb{P}(B\leq-n^{-\frac{1}{2}})$ $\lesssim\log(n)^{6}\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}n^{\frac{1}{2}}.$
Notice that we can shift the constant as a factor of the bound of $B$. As $\log(\sigma_{\min}^{-1})\sim\log(n)$, we need to choose $n$ large enough such that
$\log(n)^{6}\Big{(}\log(n)\exp(-\frac{\log(n)^{2}}{2})\Big{)}^{d}n^{\frac{1}{2}}\lesssim\frac{1}{3n}\quad\Longleftarrow\quad n\gtrsim e^{3(d+7+\frac{1}{2})},$
where the implication stems from a loose upper bound using $\log(n)^{6+d}\leq n^{6+d}$. This finishes the proof.∎
Proof of Lemma˜7.5.
From the proof of Theorem˜5.2 and as $\log(\sigma_{\min}^{-1})\sim\log(n)$ we now that
$|v_{t}(x)|\lesssim\log(n)^{2}|x|+\log(n),\quad|\tilde{v}_{t}(x)|\lesssim\log(n)^{3}.$
Thus
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$ $\lesssim\log(n)^{6}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}p_{t}(x)\;\mathrm{d}x$
$\quad+\log(n)^{4}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|x|^{2}p_{t}(x)\;\mathrm{d}x.$
By the definition of $p_{t}$ and due to the fact that the convolution of two densities is again a density, we know that
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{dy}$ $=\frac{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y\;\mathrm{d}x}$
$\leq e^{2L}\frac{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y}{\int\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y\;\mathrm{d}x}.$
Now for the inner integrals, we obtain
$\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y$ $=\exp(-\frac{|x|^{2}}{2\sigma_{t}^{2}})\int\exp(-\frac{1}{2}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}|y|^{2}+\langle\frac{t}{\sigma_{t}^{2}}x,y\rangle)\;\mathrm{d}y$
$=\exp(-\frac{|x|^{2}}{2\sigma_{t}^{2}}(2\pi)^{d/2}\big{(}\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\big{)}^{d/2}\exp(\frac{|x|^{2}}{2}\frac{t^{2}}{(t^{2}+\sigma_{t}^{2})}\sigma_{t}^{2}).$
Inserting this into the bound and collecting all factors of $|x|^{2}$, we obtain
$p_{t}(x)$ $\leq e^{2L}\frac{\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})}{\int\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})\;\mathrm{d}x}=\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2}).$
Now we bound
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}$ $p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|x|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|x|^{2}\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})\;\mathrm{d}x\;\mathrm{d}t$
$=2\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\Big{(}\int_{\log(n)}^{\infty}x_{1}^{2}\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}x_{1}^{2})\;\mathrm{d}x_{1}\Big{)}^{d}\;\mathrm{d}t$
$\leq 2\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\Big{(}\log(n)(\sigma_{t}^{2}+t^{2})\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}\log(n)^{2})\Big{)}^{d}\;\mathrm{d}t,$
where the last inequality follows from the same arguments as the bound in the proof of the first part of this Lemma. As $(\sigma_{t}^{2}+t^{2})\leq 2$, we can bound
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$ $\lesssim\log(n)^{d}\exp(-\frac{1}{4}\log^{2}(n)).$
Overall, we need to choose $n$ big enough such that
$\log(n)^{6+d}\exp(-\frac{1}{4}\log^{2}(n))\leq n^{-\frac{1}{2}}\quad\Longleftarrow\quad n\geq e^{4(\frac{1}{2}+6+d)},$
where again the implication stems from a loose upper bound using $\log(n)^{6+d}\leq n^{6+d}$.
Proof of Lemma˜7.6.
Similar to (Yakovlev and Puchkin, 2025, p.44), we bound
$\Big{|}\int_{0}^{1}\int$ $|\bar{v}_{t}^{1}(x)-v_{t}(x|y)(x)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t-\int_{0}^{1}\int|\bar{v}_{t}^{2}(x)-v_{t}(x|y)(x)|^{2}p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t\Big{|}$
$\lesssim\int_{0}^{1}\int|\bar{v}_{t}^{1}(x)-\bar{v}_{t}^{2}(x)|(|\bar{v}_{t}^{1}(x)|+|\bar{v}_{t}^{2}(x)|+2|v_{t}(x|y)|)p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t.$
For $x\in\mathbb{R}^{d}$ and $y\in[-\log(n),\log(n)]^{d}$, we have
$|v_{t}(x|y)|=\Big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x+(\sigma_{t}-t)y)\Big{|}\leq\Big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big{|}(|x|+|y|)\leq|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}|(|x|+\sqrt{d}\log(n)).$
By (33) all $\bar{v}\in\mathcal{M}$ are such that $|\bar{v}_{t}(x)|\lesssim\log(n)^{3}$. Therefore
$\int_{0}^{1}\int$ $|\bar{v}_{t}^{1}(x)-\bar{v}_{t}^{2}(x)|(|\bar{v}_{t}^{1}(x)|+|\bar{v}_{t}^{2}(x)|+2|v_{t}(x|y)|)p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t$
$\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}(\log(n)^{3}+\log(n)\int_{0}^{1}\Big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big{|}\int|x|p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t)$
$\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}\log(n)^{4}.$
The last inequality follows from Jensens inequality and the same argument as in (34). This concludes the proof. ∎
Proof of Lemma˜7.7.
For an $s$-th order partial derivative and $\{i_{1},...,i_{s}\}\subset\{1,...,d\}^{s}$, we know from the definition of $v_{t}$
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)=\mathds{1}_{s=1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(1-\frac{\sigma_{t}^{\prime}t}{\sigma_{t}})\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}\mathbb{E}_{q_{t,x}}[Y_{j}],$ (40)
where
$q_{t,x}\propto\exp(-\frac{\|x-ty\|^{2}}{2\sigma_{t}^{2}}-\frac{\|y\|^{2}}{2}-a(y))\propto\exp(\langle\eta,Y\rangle)h(y),$
where $\eta=\frac{t}{\sigma_{t}^{2}}x$ and $h(y)=\exp(-(\frac{t^{2}}{2\sigma_{t}^{2}}+\frac{1}{2})\|y\|^{2}-a(y))$. Define
$A(\eta)\coloneqq\log\int\exp(\langle\eta,Y\rangle)h(y)\;\mathrm{d}y.$
In context of exponential families, $A$ is typically called the log-partition function. Then
$\frac{\partial}{\partial\eta_{j}}A(\eta)=\mathbb{E}_{q_{t,x}}[Y_{j}].$
Thus,
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}\mathbb{E}_{q_{t,x}}[Y_{j}]=\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}\frac{\partial}{\partial\eta_{j}}A(\eta)=\big{(}\frac{t}{\sigma_{t}^{2}}\big{)}^{s}\frac{\partial^{s}}{\partial\eta_{i_{1}},...,\partial\eta_{i_{s}}}\frac{\partial}{\partial\eta_{j}}A(\eta).$
Further define the cumulant generating function as
$K(\lambda)=\log\big{(}\mathbb{E}[\exp(\langle\lambda,Y\rangle)]\big{)}.$
Then
$\mathbb{E}[\exp(\langle\lambda,Y)]=\exp(A(\lambda+\eta)-A(\eta)),\quad K(\lambda)=A(\lambda+\eta)-A(\eta).$
Differentiating in $\lambda$ yields
$\frac{\partial^{s+1}}{\partial\lambda_{i_{1}},...,\partial\lambda_{i_{s}},\partial\lambda_{j}}K(\lambda)\Big{|}_{\lambda=0}=\frac{\partial^{s+1}}{\partial\lambda_{i_{1}},...,\partial\lambda_{i_{s}},\partial\lambda_{j}}A(\lambda+\eta)-A(\eta)\Big{|}_{\lambda=0}=\frac{\partial^{s+1}}{\partial\eta_{i_{1}},...,\partial\eta_{i_{s}},\partial\eta_{j}}A(\eta).$
Hence,
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)=\mathds{1}_{s=1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(1-\frac{\sigma_{t}^{\prime}t}{\sigma_{t}})\big{(}\frac{t}{\sigma_{t}^{2}}\big{)}^{s}\kappa(Y_{i_{1}},...,Y_{i_{s}},Y_{j}),$
where $\kappa(Y_{i_{1}},...,Y_{i_{s}},Y_{j})$ is the joint cumulant, of $Y_{i_{1}},...,Y_{i_{s}},Y_{j}$, which is defined as exactly this partial derivative. For $s\geq 1$, this cumulant is shift invariant, hence
$\kappa(Y_{i_{1}},...,Y_{i_{s}},Y_{j})=\kappa(Y_{i_{1}}-\mathbb{E}[Y_{i_{1}}],...,Y_{i_{s}}-\mathbb{E}[Y_{i2}],Y_{j}-\mathbb{E}[Y_{j}]).$
Using the formula from Leonov and Shiryaev (1959) we can express the cumulant using products of mixed moments
$|\kappa(Y_{i_{1}}-\mathbb{E}[Y_{i_{1}}],...,Y_{i_{s}}-\mathbb{E}[Y_{i_{s}}],Y_{j}-\mathbb{E}[Y_{j}])|$ $=\sum_{\pi}(|\pi|-1)!\prod_{B\in\pi}\mathbb{E}\Big{(}\prod_{\ell\in B}|Y_{\ell}-\mathbb{E}[Y_{\ell}]|\Big{)}$
$\leq\big{(}\sum_{\pi}(|\pi|-1)!\big{)}\prod_{\ell=1}^{s+1}(\mathbb{E}[|Y_{B_{\ell}}-\mathbb{E}[Y_{B_{\ell}}]|^{s+1}])^{\frac{1}{s+1}},$ (41)
where $\sum_{\pi}$ represents the sum over all partitions of $\{i_{1},...,i_{s},j\}$ and $B_{\ell}$ is the $\ell$-th element of $\{i_{1},...,i_{s},j\}$.
To control the moments up to order $s+1$, we need a stronger argument than the Poincaré inequality used for the variance bound.
Definition 7.9.
A distribution $\mathbb{P}^{*}$ on $\mathbb{R}^{d}$ with density $p^{*}$ satisfies the logarithmic Sobolev inequality with $\log$-Sobolev constant $\lambda>0$ if for all smooth functions $f$
$\operatorname{Ent}(f^{2}(X))\coloneqq\mathbb{E}[f^{2}(X)\log(f^{2}(X))]-\mathbb{E}[f^{2}(X)]\log(\mathbb{E}[f^{2}(X)])\leq\frac{2}{\lambda}\mathbb{E}[|\nabla f(X)|^{2}].$ (42)
The standard Gaussian satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $\lambda=1$ (Boucheron et al., 2013, Theorem 5.4). Analogously to Poincaré inequality, its easy to see that $\mathcal{N}(\mu,\sigma^{2}I_{d})$ satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $\lambda=\sigma^{-2}$. Similar to the beginning of the proof of Theorem˜4.6, we conclude that the distribution with density (20) satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $(1+\frac{t^{2}}{\sigma_{t}^{2}})$. By (Ledoux, 2001, Lemma 1.2) we known that a bounded perturbation will again only impact the $\log$-Sobolev constant by an exponential term. Thus, we can bound the $\log$-Sobolev constant of the distribution of $Y^{x,t}$ by
$\lambda\geq e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}}).$ (43)
Now we can use Herbst’s argument (along the lines of Ledoux (1999, p.148), note that our function is so simple that we can drop the assumption of a bounded function) to show that $Y^{x,t}_{i}-\mathbb{E}[Y^{x,t}_{i}]$ is subgaussian. We omit the dependency on $x$ and $t$ in the following. Set $f(y)=e^{\frac{\vartheta y_{i}}{2}}$ for a $\vartheta>0$ in (42). Define $H(\vartheta)\coloneqq\mathbb{E}[e^{\vartheta Y_{i}}]$. Then using the dominated convergence theorem
$\operatorname{Ent}(f^{2})$ $=\mathbb{E}[\vartheta Y_{i}e^{\vartheta Y_{i}}]-\mathbb{E}[e^{\vartheta Y_{i}}]\log\big{(}\mathbb{E}[e^{\vartheta Y_{i}}]\big{)}=\vartheta H^{\prime}(\vartheta)-H(\vartheta)\log(H(\vartheta)).$
Further
$\frac{2}{\lambda}\mathbb{E}[|\nabla f(Y)|^{2}]$ $=\frac{\vartheta^{2}}{2\lambda}\mathbb{E}[e^{\vartheta Y_{i}}]=\frac{\vartheta^{2}\lambda}{2}H(\vartheta).$
Thus by (42)
$\vartheta H^{\prime}(\vartheta)-H(\vartheta)\log(H(\vartheta))\leq\frac{\vartheta^{2}}{2}H(\vartheta)\quad\Longleftrightarrow\quad\frac{1}{\vartheta}\frac{H^{\prime}(\vartheta)}{H(\vartheta)}-\frac{\log(H(\vartheta))}{\vartheta^{2}}\leq\frac{1}{2\lambda}.$
Now define $K(\vartheta)\coloneqq\frac{1}{\vartheta}\log(H(\vartheta))$. Then
$K^{\prime}(\vartheta)=\frac{H^{\prime}(\vartheta)}{H(\vartheta)}-\frac{\log(H(\vartheta))}{\vartheta^{2}}\leq\frac{1}{2\lambda}.$
As
$\lim_{\vartheta\rightarrow 0}K(\vartheta)=\lim_{\vartheta\rightarrow 0}\frac{\mathbb{E}[e^{\vartheta Y_{i}}]}{\vartheta}=\lim_{\vartheta\rightarrow 0}\frac{\mathbb{E}[Y_{i}e^{\vartheta Y_{i}}]}{\mathbb{E}[e^{\vartheta Y_{i}}]}=\lim_{\vartheta\rightarrow 0}\frac{M^{\prime}(\vartheta)}{M(\vartheta)}=\mathbb{E}[Y_{i}],$
where $M$ is the moment generating function, we can extend the definition of $K$ to
$K(\vartheta)\coloneqq\begin{cases}\frac{1}{\vartheta}\log(H(\vartheta)),&\vartheta>0,\\ \mathbb{E}[Y_{i}],&\vartheta=0.\end{cases}$
Hence
$K(\vartheta)=K(0)+\int_{0}^{\vartheta}K^{\prime}(u)\;\mathrm{d}u\leq\mathbb{E}[Y_{i}]+\frac{\vartheta}{2\lambda}.$
Going back to the function $H$, we get
$\frac{1}{\vartheta}\log(H(\vartheta))\leq\mathbb{E}[Y_{i}]+{\frac{1}{2\lambda}}\quad\Longleftrightarrow\quad H(\vartheta)\leq\exp(\vartheta\mathbb{E}[Y_{i}]+\frac{\vartheta^{2}}{2\lambda})\quad\Longleftrightarrow\quad\mathbb{E}[e^{\vartheta Y_{i}}]\leq\exp(\vartheta\mathbb{E}[Y_{i}]+\frac{\vartheta^{2}}{2\lambda}).$
Now Markov’s inequality yields
$\mathbb{P}(Y_{i}-\mathbb{E}[Y_{i}]\geq r)$ $=\mathbb{P}\big{(}\exp(\vartheta(Y_{i}-\mathbb{E}[Y_{i}]))\geq\exp(\vartheta r)\big{)}\leq\frac{\mathbb{E}[\exp(\vartheta(Y_{i}-\mathbb{E}[Y_{i}]))]}{\exp(\vartheta r)}$
$=\exp(-\vartheta r)\frac{\mathbb{E}[e^{\vartheta Y_{i}}]}{e^{\vartheta\mathbb{E}[Y_{i}]}}\leq\exp(-\vartheta r)\frac{\exp(\vartheta\mathbb{E}[Y_{i}]+\frac{\vartheta^{2}}{2\lambda})}{e^{\vartheta\mathbb{E}[Y_{i}]}}=e^{\frac{\vartheta^{2}}{2\lambda}-\vartheta r}.$
Setting $f(x)=e^{-\frac{\vartheta x_{i}}{2}}$, we obtain analogously
$\mathbb{P}(Y_{i}-\mathbb{E}[Y_{i}]\leq-r)\leq e^{\frac{\vartheta^{2}}{2\lambda}-\vartheta r}.$
Thus a union bound argument and setting $\vartheta=r\lambda$ leads to
$\mathbb{P}(|Y_{i}-\mathbb{E}[Y_{i}]|\geq r)\leq 2e^{-\frac{r^{2}\lambda}{2}}.$
Now we can bound the $s+1$-th moment along the lines of (Vershynin, 2018, Proposition 2.5.2) via
$\mathbb{E}[|Y_{i}-\mathbb{E}[Y_{i}]|^{s+1}]$ $=\int_{0}^{\infty}\mathbb{P}(|Y_{i}-\mathbb{E}[Y_{i}]|^{s+1}\geq u)\;\mathrm{d}u=\int_{0}^{\infty}\mathbb{P}(|Y_{i}-\mathbb{E}[Y_{i}]|\geq r)r^{s}(s+1)\;\mathrm{d}r$
$\leq 2(s+1)\int_{0}^{\infty}e^{-\frac{r^{2}\lambda}{2}}r^{s}\;\mathrm{d}r=2^{\frac{s+1}{2}}\lambda^{-\frac{s+1}{2}}\Gamma(\frac{s+1}{2}).$
Taking the $s+1$-th root leads to the bound
$\mathbb{E}[|Y_{i}-\mathbb{E}[Y_{i}]|^{s+1}]^{\frac{1}{s+1}}$ $\lesssim\lambda^{-\frac{1}{2}}.$
Plugging in the $\log$ Sobolev constant of $Y$ from (43), we can control all moments via
$\mathbb{E}[|Y_{i_{j}}^{x,t}-\mathbb{E}[Y_{i_{j}}^{x,t}]|^{s+1}]^{\frac{1}{s+1}}$ $\lesssim\frac{\sigma_{t}}{\sqrt{\sigma_{t}^{2}+t^{2}}}\leq\min\big{(}1,\frac{\sigma_{t}}{t}\big{)}.$
Thus we can bound (41) by
$\prod_{\ell=1}^{s+1}(\mathbb{E}[|Y_{B_{\ell}}-\mathbb{E}[Y_{B_{\ell}}]|^{s+1}])^{\frac{1}{s+1}}\lesssim\min\Big{(}1,\frac{\sigma_{t}^{s+1}}{t^{s+1}}\Big{)}.$
We obtain for
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)\lesssim\mathds{1}_{s=1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(1-\frac{\sigma_{t}^{\prime}t}{\sigma_{t}})\big{(}\frac{t}{\sigma_{t}^{2}}\big{)}^{s}\min\Big{(}1,\frac{\sigma_{t}^{s+1}}{t^{s+1}}\Big{)}.$
Minimizing in $\sigma_{t}$ and using $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min})$ leads to
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)\lesssim\log(\sigma_{\min})\sigma_{\min}^{-s+1}.$
Proof of Lemma˜7.8.
Define
$q_{x,t}(y)\coloneqq\frac{\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))}{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y},\quad w_{x,t}(y)\coloneqq\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y)).$
For $k=1$ we obtain using the dominated convergence theorem, the quotient rule and the derivative of the logarithm
$\frac{\partial}{\partial t}v^{j}_{t}(x)$ $=\frac{\partial}{\partial t}\mathbb{E}_{Y\sim q_{x,t}}[v_{t}(x|Y)_{j}]$ (44)
$=\mathbb{E}_{Y\sim q_{x,t}}\Big{[}\frac{\partial}{\partial t}v_{t}(x|Y)_{j}\Big{]}+\mathbb{E}_{Y\sim q_{x,t}}\Big{[}v_{t}(x|Y)_{j}\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big{]}$
$\quad-\mathbb{E}_{Y\sim q_{x,t}}\Big{[}v_{t}(x|Y)_{j}\Big{]}\mathbb{E}_{Y\sim q_{x,t}}\Big{[}\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big{]}$
$=\mathbb{E}_{Y\sim q_{x,t}}\Big{[}\frac{\partial}{\partial t}v_{t}(x|Y)_{j}\Big{]}+\operatorname{Cov}\Big{(}v_{t}(x|Y)_{j},\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big{)}.$
We can use the same structure to obtain higher derivatives. Thus, all terms occurring in an derivative of $v_{t}$ of order $k$ are either derivatives of $v_{t}(x|y)_{j}$, derivatives of $\log(w_{x,t}(Y))$ (including the functions themselves as 0-th order derivative) or products of these derivatives. Since
$v_{t}(x|y)_{j}$ $=\log(\sigma_{\min})(x_{j}-ty_{j})+y_{j},$ (45)
$\frac{\partial}{\partial t}v_{t}(x|y)_{j}$ $=\log(\sigma_{\min})y_{j},$
all higher derivatives of $v_{t}(\cdot|\cdot)$ vanish. For the second term occurring in (44), we have that
$\frac{\partial}{\partial t}\log(w_{x,t}(Y))$ $=-\frac{\partial}{\partial t}\frac{1}{2}(|x-ty|^{2})\sigma_{t}^{-2}$
$=(\langle x,y\rangle-t|y|^{2})\sigma_{t}^{-2}+\frac{1}{2}|x-ty|^{2}\cdot 2\log(\sigma_{\min})\sigma_{t}^{-2}.$ (46)
We conclude that
$\frac{\partial^{k}}{\partial t^{k}}\log(w_{x,t}(Y))$ $\lesssim\operatorname{poly}(x,y,t)\operatorname{polylog}(\sigma_{\min}^{-1})\sigma_{t}^{-2},$ (47)
where $\operatorname{poly}(x,y,t)$ is a again polynomial of finite degree in the components of $x,y,t$. Thus, bounds that are worse than $\sigma_{t}^{-2}$ can only appear when $\frac{\partial}{\partial t}\log(w_{x,t}(Y))$ occurs raised to a power. Looking at (44), we see that the highest power of $\frac{\partial}{\partial t}\log(w_{x,t}(Y))$ increases by $1$ each time the derivative is taken via the covariance term. Hence in the $k$-th derivative, the highest order terms are of the form
$\operatorname{Cov}(f(Y)(\frac{\partial}{\partial t}\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y))),$
where either $f\equiv 1$ or $f$ is of the form (45) or (47). First, let $f\equiv 1$. Then due to (46)
$\operatorname{Cov}((\frac{\partial}{\partial t}$ $\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y)))$
$=\mathbb{E}\Big{[}\Big{(}\frac{\partial}{\partial t}\log(w_{x,t}(Y))-\mathbb{E}\Big{[}\frac{\partial}{\partial t}\log(w_{x,t}(Y))\big{]}\Big{)}^{k}\Big{]}$
$=\sigma_{t}^{-2k}\mathbb{E}\Big{[}\Big{(}\langle(1-t2\log(\sigma_{\min}))x,Y\rangle-\mathbb{E}[\langle(1-t2\log(\sigma_{\min}))x,Y\rangle]$
$\qquad\qquad+(t^{2}2\log(\sigma_{\min})-t)|Y|^{2}-\mathbb{E}[(t^{2}2\log(\sigma_{\min})-t)|Y^{2}|]\Big{)}^{k}\Big{]}$
$\lesssim\sigma_{t}^{-2k}\sum_{i=1}^{d}(1-t2\log(\sigma_{\min}))^{k}x^{k}\mathbb{E}\Big{[}\big{(}Y_{i}-\mathbb{E}[Y_{i}]\big{)}^{k}\Big{]}+\sum_{i=1}^{d}(t^{2}2\log(\sigma_{\min})t-t)^{k}\mathbb{E}\Big{[}\big{(}Y_{i}^{2}-\mathbb{E}[Y_{i}^{2}]\big{)}^{k}\Big{]},$
where all expectations are taken with respect to $Y\sim q_{x,t}$. From the proof of Lemma˜7.7 we know that
$\mathbb{E}\Big{[}\big{(}Y_{i}-\mathbb{E}[Y_{i}]\big{)}^{k}\Big{]}\lesssim\big{(}\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\big{)}^{\frac{k}{2}}\leq\min(\frac{\sigma_{t}^{2}}{t^{2}},1)^{\frac{k}{2}}=\min(\frac{\sigma_{t}^{k}}{t^{k}},1).$
By Vershynin (2018, Lemma 2.7.6) combined with Vershynin (2018, Lemma 2.7.1) we know that if $Y_{i}-\mathbb{E}[Y_{i}]$ is subgaussian, which was shown in Lemma˜7.7, then $(Y_{i}-\mathbb{E}[Y_{i}])^{2}$ is subexponential with the same order of subgaussian and subexponential norm. For a definition of a subexponential random variables, the subexponential norm and the subgaussian norm, we refer to Vershynin (2018, Section 2.7). Note that all the constants in Vershynin (2018, Proposition 2.7.1) only differ by a constant.
$\mathbb{E}\Big{[}\big{(}Y_{i}^{2}-\mathbb{E}[Y_{i}^{2}]\big{)}^{k}\Big{]}$ $=\mathbb{E}\Big{[}\big{(}(Y_{i}-\mathbb{E}[Y_{i}])^{2}-\mathbb{E}[(Y_{i}-\mathbb{E}[Y_{i}])^{2}]+2\mathbb{E}[Y_{i}](Y_{i}-\mathbb{E}[Y_{i}])\big{)}^{k}\Big{]}$
Note that $\mathbb{E}[Y_{i}]\lesssim\operatorname{polylog}(n)$ as $x\in[-\log(n),\log(n)]^{d}$. Hence $Y_{i}^{2}-\mathbb{E}[Y_{i}^{2}]$ can be represented by the sum of a subexponential and a subgaussian variable. Furthermore, if $Y_{i}-\mathbb{E}[Y_{i}]$ is subgaussian then it is also subexponential and the subexponential norm differs only by a constant from the subgaussian norm. Additionally, from Vershynin (2018, Definition 2.7.5) and the properties of a norm we learn that the sum of subexponential random variables is again subexponential and that we can bound the subexponential norm of the sum by the sum of the individual subexponential norms. Using that the subgaussian norm of $Y_{i}-\mathbb{E}[Y_{i}]$ is of order $\frac{\sigma_{t}}{\sqrt{\sigma_{t}^{2}+t^{2}}}$, we obtain using Vershynin (2018, Proposition 2.5.2)
$\mathbb{E}\Big{[}\big{(}Y_{i}^{2}-\mathbb{E}[Y_{i}^{2}]\big{)}^{k}\Big{]}$ $\lesssim\big{(}\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\big{)}^{\frac{k}{2}}\leq\min(\frac{\sigma_{t}^{k}}{t^{k}},1).$
Now we consider the case that $f$ is of the form (45) or (47). First, we note that as shown in the proof of Theorem˜4.6, the distribution with density $q_{x,t}$ satisfies the Poincaré inequality with a constant
$\rho\geq e^{-2L}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}.$
An immediate consequence of the Poincaré inequality Definition˜7.3 is the following bound for all smooth functions $g,h$
$\operatorname{Cov}(g(Y),h(Y))\leq\frac{1}{\rho}\sqrt{\mathbb{E}[|\nabla g(Y)|^{2}]\mathbb{E}[|\nabla h(Y)|^{2}]}.$ (48)
We are going to use this to bound
$\operatorname{Cov}(f(Y)(\frac{\partial}{\partial t}\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y))).$
Define
$g(y)\coloneqq f(y)(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-1}.$
Then
$\nabla g(y)=(\nabla f(y))(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-1}+f(y)(k-1)\nabla(\frac{\partial}{\partial t}\log(w_{x,t}(y)))^{k-2}(\nabla\frac{\partial}{\partial t}\log(w_{x,t}(y)))$
and
$|\nabla g(y)|^{2}\lesssim|\nabla f(y)|^{2}|\frac{\partial}{\partial t}\log(w_{x,t}(y)))|^{2(k-1)}+|f(y)|^{2}(k-1)^{2}|(\frac{\partial}{\partial t}\log(w_{x,t}(y)))|^{2(k-2)}|\nabla\frac{\partial}{\partial t}\log(w_{x,t}(y))|^{2}$
$\lesssim|\nabla f(y)|^{2}\big{|}\sigma_{t}^{-4(k-1))}|x-ty|^{2(k-1)}|y|^{2(k-1)}\big{|}+\big{|}\sigma_{t}^{-4(k-1)}\frac{1}{2}|x-ty|^{4(k-1)}(2\log(\sigma_{\min}))^{2(k-1)}\big{|}$ (49)
$\quad+|f(y)|^{2}\big{(}\big{|}\sigma_{t}^{-4(k-2)}|x-ty|^{2(k-2)}|y|^{2(k-2)}\big{|}+\big{|}\sigma_{t}^{-4(k-2)}\frac{1}{2^{2(k-1)}}|x-ty|^{4(k-2)}(2\log(\sigma_{\min}))^{2(k-2)}\big{|}\big{)}$ (50)
$\qquad\cdot|\sigma_{t}^{-2}(x-2ty+t(x-ty)2\log(\sigma_{\min}))|^{2}.$ (51)
All possible functions $f$ are polynomials in $y$, hence the derivatives are also polynomials. As $v_{t}(\cdot|\cdot)$ is linear in $y_{i}$ and $\frac{\partial}{\partial t}\log(w_{x,t})$ is quadratic in the components, the derivatives are of that degree or lower. We denote this by $\operatorname{poly}_{2}(x,y,t)$. In case $f$ is of the form (47) $\nabla f$ and $f$ are of order $\sigma_{t}^{-2}$, else the dependency on $\sigma_{\min}$ or $\sigma_{t}$ is only logarithmically. For the first term in (49)
$\mathbb{E}$ $[|\nabla f(Y)|^{2}\sigma_{t}^{-4(k-1)}|x-tY|^{2(k-1)}|Y|^{2(k-1)}\big{]}$
$=\int\operatorname{polylog}(\sigma_{\min})\operatorname{poly}_{2}(x,y,t)\sigma_{t}^{-4(k-1))}|x-ty|^{2(k-1)}|y|^{2(k-1)}\frac{\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))}{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}\;\mathrm{d}y$
$\leq\frac{e^{2L}\operatorname{polylog}(\sigma_{\min})}{\int\exp(-\frac{|z|^{2}}{2}-\frac{|\frac{x-\sigma_{t}z}{t}|^{2}}{2})\;\mathrm{d}y}$
$\quad\cdot\int\operatorname{poly}_{2}(x,\frac{x-\sigma z}{t},t)\sigma_{t}^{-4(k-1)}|\sigma_{t}z|^{2(k-1)}|\frac{x-\sigma_{t}z}{t}|^{2(k-1)}\exp(-\frac{|z|^{2}}{2}-\frac{|\frac{x-\sigma_{t}z}{t}|^{2}}{2})\;\mathrm{d}y$
$\lesssim\operatorname{polylog}(\sigma_{\min})\sigma_{t}^{-2k-2}|x|^{2(k-1)},$
where the last inequality stems from similar arguments as in the proof of Theorem˜4.6 and Theorem˜5.2 for higher order moments. The exact calculations are omitted for the sake of brevity. The same calculation for the other terms in (49), (50), (51) yield
$\mathbb{E}[|\nabla g(y)|^{2}]\lesssim\operatorname{polylog}(\sigma_{\min})\log(n)^{4(k-1)}\sigma_{t}^{-2k-4},$
where we used that $x\in[-\log(n),\log(n)]^{d}$ to combine the worst case dependencies. For $h$, we obtain in similar manner
$\mathbb{E}[|\nabla h(y)|^{2}]\lesssim\sigma_{t}^{-4}\log(n)^{4}.$
Inserting these bounds in (48), we obtain
$\operatorname{Cov}(f(Y)(\frac{\partial}{\partial t}\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y)))$ $\lesssim\operatorname{polylog}(\sigma_{\min})\operatorname{polylog}(n)\frac{\sigma_{t}^{2}}{\sigma_{t}^{2}+t^{2}}\sigma_{t}^{-k-4}.$
Hence in all cases, the terms occurring in the $k$-th derivative are bounded by
$\frac{\partial^{k}}{\partial t^{k}}v^{j}_{t}(x)\lesssim\operatorname{polylog}(\sigma_{\min})\operatorname{polylog}(n)\sigma_{\min}^{-k-2}.$
This concludes the proof. ∎
7.4.1 Helper results
Lemma 7.10.
Let $\sigma_{t}=\sigma_{\min}^{t}$ with $\sigma_{\min}\in(0,1)$. Then for $t\in(0,1)$
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\max(\log(\sigma_{\min}^{-1}),e^{2}).$
Proof.
First, let $t\geq\frac{1}{\log(\sigma_{\min}^{-1})}$. Then
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\frac{1}{t}\leq\log(\sigma_{\min}^{-1}).$
If $t<\frac{1}{\log(\sigma_{\min}^{-1})}$, then
$\sigma_{\min}^{2t}=\exp(2\log(\sigma_{\min})t)\geq\exp(-2)$
and thus
$\frac{t}{t^{2}+\sigma_{t}^{2}}\leq\frac{t}{t^{2}+\exp(-2)}\leq\exp(2).$
Lemma 7.11.
For $p^{*}$ of the form (14) and any Lipschitz $1$ function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$,
$\Big{|}\int\nabla f(z)p^{*}(x-z)\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big{|}\lesssim|x-y|.$ (52)
Proof.
Using Hölders inequality we obtain
$\Big{|}\int\nabla f(z)p^{*}(x-z)\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big{|}$ $\leq\|\nabla f\|_{\infty}\int|p^{*}(x-z)-p^{*}(y-z)|\;\mathrm{d}z$
$\leq\sqrt{d}\int\Big{|}\int_{0}^{1}\langle x-y,\nabla p^{*}(x+t(y-x)-z)\rangle\;\mathrm{d}t\Big{|}\;\mathrm{d}z,$
where the last inequality follows from the Lipschitz bound on $f$ and a backwards-application of the multivariate chain rule. The $\nabla$-operator is used with respect to the $\mathbb{R}^{d}$ valued input of $p^{*}$. Using the Cauchy-Schwarz inequality and changing the order of integration, we conclude
$\int\Big{|}\int_{0}^{1}\langle x-y,\nabla p^{*}(x+t(y-x)-z)\rangle\Big{|}\;\mathrm{d}z$ $\leq|x-y|\int\int_{0}^{1}|\nabla p^{*}(x+t(y-x)-z)|\;\mathrm{d}t\;\mathrm{d}z$
$=|x-y|\int_{0}^{1}\int|\nabla p^{*}(w)|\;\mathrm{d}w\;\mathrm{d}t.$
We can bound the integral over the derivative via
$\int|\nabla p^{*}(w)|\;\mathrm{d}w$ $=\int|-w-\nabla a(w)|\frac{\exp(-\frac{-|w|^{2}}{2}-a(w))}{\int\exp(-\frac{-|w|^{2}}{2}-a(w))\;\mathrm{d}w}\;\mathrm{d}w$
$\leq\mathbb{E}_{W\sim p^{*}}[|W|]+\sqrt{d}L.$
We can further bound
$\mathbb{E}_{W\sim p^{*}}[|W|]$ $\leq e^{2L}\mathbb{E}_{W\sim\mathcal{N}(0,I_{d})}[|X|]\leq e^{2L}\sqrt{\mathbb{E}_{W\sim\mathcal{N}(0,I_{d})}[|X|^{2}]}=e^{2L}\sqrt{d}.$
Thus
$\Big{|}\int\nabla f(z)p^{*}(x-z)\mathrm{d}z-\int\nabla f(z)p^{*}(y-z)\;\mathrm{d}z\Big{|}\leq d(e^{2L}+L)|x-y|.$